### v1

# LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis

Abstract
* Equal contribution. Correspondence to Ruibo Fu: ruibo.fu@nlpr.ia.ac.cn.
Portrait image animation using audio has rapidly advanced, enabling the creation of increasingly realistic and expressive animated faces. The challenges of this multimodality-guided video generation task involve fusing various modalities while ensuring consistency in timing and portrait. We further seek to produce vivid talking heads. To address these challenges, we present LetsTalk (LatEnt Diffusion TranSformer for Talking Video Synthesis), a diffusion transformer that incorporates modular temporal and spatial attention mechanisms to merge multimodality and enhance spatial-temporal consistency. To handle multimodal conditions, we first summarize three fusion schemes, ranging from shallow to deep fusion compactness, and thoroughly explore their impact and applicability. Then we propose a suitable solution according to the modality differences of image, audio, and video generation. For portrait, we utilize a deep fusion scheme (Symbiotic Fusion) to ensure portrait consistency. For audio, we implement a shallow fusion scheme (Direct Fusion) to achieve audio-animation alignment while preserving diversity. Our extensive experiments demonstrate that our approach generates temporally coherent and realistic videos with enhanced diversity and liveliness. The project page is available at Let’s Talk.
{strip}
[Uncaptioned image][Uncaptioned image]
Figure 1: We propose LetsTalk, a diffusion-based transformer for audio-driven portrait image animation. Left: Given a single reference image and audio, LetsTalks can produce a realistic and vivid video aligned with the input audio. Note that each column corresponds to the same audio. The results show that LetsTalk can drive consistent and reasonable mouth motions for the input audio. Right: Generation quality vs. inference time on the HDTF dataset, the circle area reflects the number of parameters of the method. Our LetsTalk achieves the best quality while also being highly efficient in inference, compared to current mainstream diffusion-based methods such as Hallo and AniPortrait. In addition, our base version (LetsTalk-B) achieves performance similar to Hallo with only 8 $\times$ fewer parameters.
1 Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. However, producing reasonable animation faces the following main challenges: consistency in timing and portrait, audio-animation alignment, and diversity.
To overcome these challenges, several methods [52, 45] adopt a 3D morphable model (3DMM) [4, 26, 3] to represent the reference portrait and use audio signals to drive the 3DMM. They feed the 3D signals (e.g. coefficients, poses) from the driven 3DMM into a neural renderer to generate a talking head. While achieving great consistency in timing and portrait, they tend to produce dull results, such as unblinking eyes and fixed hair, which damage the diversity of animations. In contrast to these 3D-aware methods, some other methods [7, 32, 55, 10, 54] mainly focus on capturing the correspondence between the given audio and related 2D movements, like facial muscle and lip motions, directly aligning animation with audio. They implicitly learn the correspondence between audio and animation in the latent space, animating the portrait image by feeding forward the audio. However, these methods still struggle to disentangle the components representing the static and dynamic aspects of the face, even hindering the temporal consistency and diversity of videos.
Recently, diffusion models [18, 13, 34] have emerged as a promising technique to generate realistic 2D images [34, 51] and have even achieved remarkable results in the video generation [29, 28]. Notably, AnimateAnyone [21] incorporates motion modules and ReferenceNet into diffusion models to generate realistic and pose-controllable human animation, which verifies that the end-to-end diffusion models can seamlessly capture the correspondence between the condition signal and human animation. Inspired by it, recent methods [47, 9, 11] replace the pose signals in AnimateAnyone with audio signals and achieve talking head generation. Leveraging diffusion models, these methods excel at generating various unconscious actions, thereby enhancing diversity. While achieving promising results, these methods adopt U-Net as the backbone and lack the flexibility for better fusing multimodality. In addition, although the U-Net backbone works well for image generation, it introduces considerable payloads and struggles with scalability for video generation. Therefore, it is necessary to propose a transformer framework to solve this task and reasonably fuse multimodality, ensuring better portrait consistency and producing more vivid characters.
In this work, we present LetsTalk, a latent diffusion transformer to solve this video generation-like task via the diffusion paradigm. In LetsTalk, we first employ a variational autoencoder (VAE) to reduce this task into a generation task within low-resolution latent space. To enhance the spatial-temporal consistency, we equip each transformer block with temporal and spatial attention modules to capture intra-frame and cross-frame coherence. Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference image and audio sequences. To better exploit multimodality to guide video generation, we summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, which uses a fixed modality to guide all features in the backbone through a cross-attention mechanism; 2) Siamese Fusion, like ReferenceNet, which feeds modality into a backbone-like network and uses the corresponding features to guide the features in the backbone; 3) Symbiotic Fusion, which concatenates modality with the input at the beginning and achieves guidance only through backbone’s inherent self-attention mechanisms. After summarizing these fusion schemes, we explore the impact of adopting different schemes for the given reference image and audio sequence. To better guide video generation, we use Symbiotic Fusion for the reference image whose modality is close to the video while strongly ensuring the portrait consistency of the generated video. Given the audio, we ask the animation to be aligned with it, generating plausible animation, especially lip motions, and further pursuing diversity (e.g. blinking, head moving, and expressions). Considering the audio modality differs from video, we adopt the Direct Fusion that ensures the audio-animation alignments and avoids ambiguity caused by significant modality differences in the fusion. Additionally, we find that unconscious portrait actions, closely related to diversity, are primarily generated during the denoising of diffusion models, and the shallow fusion scheme effectively preserves the diversity. The empirical results indicate that our innovative framework realizes realistic and high-fidelity talking-head synthesis, advancing the frontiers of realism and expressiveness in audio-driven video generation. To the best of our knowledge, our method is the first work to introduce the diffusion transformer for audio-driven portrait animation.
We summarize the contributions of this work as follows:
  • We present LetsTalk that pioneers the utilization of transformers in diffusion-based, audio-driven portrait animation by introducing modular spatial-temporal attention mechanisms;
  • We summarize and deeply analyze various multimodal fusion schemes, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality;
  • Our experiments on standard benchmarks demonstrate the superiority of our methods to other approaches in driving portrait animation via audio.
2 Related Work
2.1 Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [18, 13] and video generation [17, 21, 20, 57]. A notable example is Stable Diffusion [34], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [37]. This approach has been extended to more complex tasks like video generation [17, 21], where models such as Video Diffusion Models (VDM)[20] apply the diffusion process with space-time factorized training, and ImagenVideo[19] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [43], AnimateDiff [17], and VideoCrafter [6], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [31], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [29] and VDT [28], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2 Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [32] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [7] and MakeItTalk [55] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [52], Genface [50], and DiffTalk [38] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [39], DreamTalk [30], and VividTalk [40], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [48], EMO [41] and AniPortrait [45] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [27] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [16] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [47, 11] and X-Portrait [46] have adopted the latent diffusion framework pioneered in AnimateAnyone [21] or ControlNet [51], adding more control over elements such as lip movements, expressions, and poses.
2.3 Conditionally Guided Diffusion Model
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [33, 1, 5, 49, 24]. ControlNet [51] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [21] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [15] integrates structural guidance through depth maps and cross-attention, while MCDiff [8], LaMD [22], and MotionCtrl [44] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [38, 39]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [41, 47]. While control over both images and videos in latent diffusion models has seen significant advancements, diffusion transformers remain underdeveloped in this area, particularly when it comes to handling multiple modalities (e.g. image, video, audio). To address this gap, we explore various diffusion conditioning strategies tailored to the transformer architecture, aiming to achieve better control and enhanced generation quality.
(a) Overview of LetsTalk
(b) Transformer Block
Figure 2: The overview of our method (a) and the illustration of our designed transformer block (b). For better illustration, we omit the timestep encoder and Layer Norm in (b). LetsTalk integrates transformer blocks equipped with both temporal and spatial attention modules, designed to capture intra-frame spatial details and establish temporal correspondence across time steps. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding.
3 Method
We present LetsTalk, a diffusion transformer for portrait image animation driven by speech, as shown in Fig. 2. In this section, we first briefly introduce diffusion models and the formulation of our task in Sec. 3.1. Then, we delve into the details of our transformer backbone in Sec. 3.2. Finally, we summarize multimodal fusion schemes and present our adopted strategy for fusing the input reference portrait and audio in Sec. 3.3.
3.1 Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,1)$, forming a Markov chain over $T$ steps.
Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$. This is optimized via the variational lower bound of $\log p(x_{0})$:
$\begin{split}\vspace{-0.3cm}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances. Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [25] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
Since we follow DiT and introduce a pretrained VAE in our framework, our goal is reduced from generating a video $\bm{V}\in\mathbb{R}^{F\times\hat{H}\times\hat{W}\times 3}$ to generating a latent feature $\bm{F}\in\mathbb{R}^{F\times H\times W\times C}$, where $F$ denotes the number of video frames333 The encoder in VAE downsamples the image by a factor of $8$, and $\hat{H}=8\times H,\hat{W}=8\times W$ We first separate the latent feature $\bm{F}$ into non-overlapping square patches along the spatial dimensions (i.e., $H$ and $W$) and obtain a patch sequence of length $F\times P$. For each patch, we flatten it and use an MLP to project it into the latent space as $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, where $d$ is the dimension of the latent space. Then we add spatial and temporal positional embeddings [42] to the latent feature and make it sensitive to spatial and temporal information. Finally, we feed the processed latent feature into transformers and perform denoising.
(a) Direct Fusion
(b) Siamese Fusion
(c) Symbiotic Fusion
Figure 3: Illustration of three multimodal fusion schemes, our transformer backbone is formed by the left-side blocks. (a) Direct Fusion. Directly feeding condition into each block’s cross-attention module; (b) Siamese Fusion. Maintaining a similar transformer and feeding the condition into it, extracting the corresponding features to guide the features in the backbone; (c) Symbiotic Fusion. Concatenating modality with the input at the beginning, then feeding it into the backbone, achieving fusion via the inherent self-attention mechanisms.
3.2 Efficient Spatial-Temporal-Aware Transformer Backbone
Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we feed it into a transformer and predict the noise to be removed. For video generation, we extend the backbone in DiT that only consists of spatial transformer blocks. It is straightforward to flatten the $\bm{Z}$ into $F\times P$ tokens with shape $(F\times P,d)$ and repeatedly conduct a self-attention mechanism to achieve spatiotemporal fusion. However, this approach results in a huge payload for computing the attention matrix with shape $(F\times P,F\times P)$, as it grows with the square of the increase in resolution or frames.
To address this problem, we propose to achieve spatiotemporal fusion in an “interleaved fusion” manner, implemented by alternately stacking spatial and temporal blocks. Considering that $F$ refers to the temporal dimension (i.e. the number of frames) and $P$ refers to the spatial dimension (i.e. the number of patches in each frame), to feed the feature into the spatial block, we maintain the feature with shape $(F,P,d)$ and calculate the attention matrix with shape $(F,P,P)$ to capture spatial information exclusively among patches in the same frame. Similarly, to feed the feature into the temporal block, we maintain the feature with shape $(P,F,d)$ and calculate the attention matrix with shape $(P,F,F)$ to capture temporal information among frames in the same patch. Compared with the straightforward scheme that produces an attention matrix with $F^{2}\times P^{2}$ parameters, our scheme significantly reduces the amount to $F\times P^{2}+P\times F^{2}$. Our experiments verify the efficacy of our backbone extension in terms of image quality and temporal consistency.
3.3 Multimodal Fusion Schemes for Portrait and Speech Audio
After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task. Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
  • Direct Fusion (Fig. 3(a)) directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention.
  • Siamese Fusion (Fig. 3(b)) feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
  • Symbiotic Fusion (Fig. 3(c)) concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
After defining these schemes with fusion compactness ranging from shallow to deep, we aim to adopt reasonable solutions for input modalities (i.e. the reference image of portrait and audio).
Given the reference image, we require that the character of the generated video is consistent with the reference, i.e. using the reference image to strongly guide video generation. Moreover, Since the image modality is close to the video, Symbiotic Fusion is appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ (i.e. frames) times and concatenate it with the latent feature $\bm{F}$ along the spatial dimension [^4] To concatenate latent feature $\bm{F}\in\mathbb{R}^{F\times P\times d}$ with the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$, we extend and repeat the portrait feature along the temporal dimension as $\tilde{\bm{P}}\in\mathbb{R}^{F\times P\times d}$, the shape of concatenated latent feature is $[F,2P,d]$. and feed them into the transformer backbone. In contrast, we observe that Direct Fusion attempts to use shallow portrait feature $\bm{P}$ to guide the learning of all blocks, leading to misaligning the features in the deeper blocks. While Siamese Fusion addresses this misalignment and achieves comparable results to Symbiotic Fusion, it needs to maintain a siamese transformer similar to the backbone, which introduces additional model parameters and complicates the learning task.
Datasets Raw Filtered
#IDs #Hours #IDs #Hours
HDTF 411 15.75 356 13.95
CelebV-HQ 35,666 68 22,423 41.27
Train - - 20,501 49.69
Eval - - 2,278 5.522
Figure 4: The dataset splits for HDTF and CelebV-HQ, including raw/filtered data and training/evaluation distributions.
As for the audio, we require it to reasonably drive the animation, especially the lip motion [52]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Therefore, we aim for the audio to have a minimal correlation with these animations. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [47] and extract richer semantic information from different layers of the wav2vec encoder [36]. A sliding window mechanism is employed to capture long-term sequential audio information. Through nonlinear mapping, we obtain audio tokens enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
4 Experiment
4.1 Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [53] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [47]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 22,423 clips with 41.27 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [23]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [12] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [30], SadTalker (138M) [52], AniPortrait (2.8B) [45], and Hallo (2.4B) [47]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes, as shown in Fig. 4.
Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 105.434 529.510 0.406 3.466
SadTalker 102.371 362.634 0.414 2.168
AniPortrait 78.284 331.117 0.623 2.903
Hallo 45.754 285.988 0.698 2.235
Ours (Base) 46.240 289.385 0.666 2.587
Ours (Large) 37.134 219.119 0.632 1.962
Table 1: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 64.297 293.905 0.448 5.915
SadTalker 42.017 77.802 0.469 3.372
AniPortrait 33.408 107.236 0.578 3.722
Hallo 8.504 49.431 0.580 3.280
Ours (Base) 9.585 36.159 0.598 3.349
Ours (Large) 7.610 33.175 0.584 3.038
Table 2: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
Implementation details Our approach differs from previous methods [41, 47, 45] that require a two-stage training process. We adopt an end-to-end training strategy. Thanks to the scalability of transformers, we set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability. Following common practices in generative modeling works [31, 2], an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Please refer to Sec. 7 for more implementation details.
4.2 Quantitative Results
The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab. 1 and Tab. 2, respectively. The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment. Notably, our approach achieves the lowest FID (37.134) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset. These results show that our method generates both realistic and temporally consistent talking-head videos. For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
4.3 Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 9) and the CelebV-HQ dataset (Fig. 10). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment. More results can be found in Sec. 8.
4.4 Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 3 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training. We conduct experiments on the HDTF dataset using the LetsTalk-B model to validate the impact of different fusion schemes.
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab. 3. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [52] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 4 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation. To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^5] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Figure 7: Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Direct Fusion 52.999 389.01 0.497 3.149
Siamese Fusion 47.779 286.692 0.651 2.597
Symbiotic Fusion 46.240 289.385 0.666 2.587
Table 3: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Symbiotic Fusion 51.880 403.628 0.471 3.398
Siamese Fusion 48.041 318.637 0.585 2.706
Direct Fusion 46.240 289.385 0.666 2.587
Table 4: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
5 Conclusion
We present LetsTalk, a novel multimodality-guided diffusion transformer to achieve superior audio-driven portrait animation results using audio input. In LetsTalk, we introduce modular spatial-temporal attention mechanisms into our transformer backbone to enhance temporal consistency. We summarize three fusion schemes to handle multimodality and deeply analyze their impact. We then adopt the optimal fusion schemes for input reference portrait and audio, respectively. Our extensive experiments demonstrate the effectiveness of LetsTalk in achieving state-of-the-art portrait animation results with a good balance between performance and payload.
Supplementary Material
6 Network Architecture
6.1 Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) [25] from Stable Diffusion [34]. Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 5. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
In the spatial-temporal-aware block, we integrate the block details into our pipeline to model the relationship between temporal and spatial attention. Inspired by  [28], we introduce a fully connected layer between the temporal and spatial attention mechanisms. This layer applies a linear transformation to the output of the temporal attention, enhancing the learning of temporal features and ensuring consistent dimensionality for subsequent processing.
Model Patch size Layer num N Hidden size D Paramaters
LetsTalk-B 2 12 768 193M
LetsTalk-L 2 24 1024 663M
Table 5: Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
6.2 Audio Encoder
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec [36], employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
6.3 Siamese Transformer
The Siamese Fusion scheme has been adopted in AnimateAnyone [21], EMO [41], Hallo [47], and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT [14] pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
7 Additional Implementation Details
7.1 Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM [13], specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
7.2 Long-Duration Generation
Figure 8: The illustration of the long-duration generation process.
Unlike the previous frame-by-frame generation approach [38, 39], our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we follow [47] and pick the last $N$ frames from the last generated clip as the initial frames of the current clip ($N=2$ in our experiments). The initial frames are solely involved in guiding the temporal attention, ensuring continuity between consecutive clips during sampling. Fig. 8 illustrates the long-duration generation process.
7.3 Training
We initialize LetsTalk from a pre-trained Latte [29] model on the FaceForensics dataset [35]. Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
7.4 Inference
During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio. The length and resolution of the generated video clip during inference need to remain consistent with those used during training, which in our experiments are $16$ frames and a resolution of $256\times 256$. We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
8 Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF [53] and Celeb-HQ [56] benchmarks ( Fig. 9 and Fig. 10).
We show the portrait animation results of LetsTalk driven by English (Fig. 11), Chinese (Fig. 12), and singing audios (Fig. 13). We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig. 14). Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
Figure 9: The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 10: The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
9 Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF [53] and CelebV-HQ [56], which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^6] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 11: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 12: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 14: Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.
10 Limitation and Future Work
Despite the advancements proposed in this study regarding portrait image animation, several limitations necessitate further exploration and consideration. (1) Resource and Dataset Limitations: Unlike recent works such as Hallo [47] and EMO [41], which use diverse real-world datasets, our study is limited by the availability of publicly accessible datasets. These datasets are relatively small and may not be as robust, thus limiting the generalizability of LetsTalk. Additionally, the computational resources available were insufficient to fully exploit larger models, restricting us to smaller-scale models with fewer parameters. Future research could benefit from larger datasets and scaling up model sizes to explore the scalability of LetsTalk. (2) Long-Duration Animation: A key limitation in generating long-duration animations with our end-to-end framework is the challenge of maintaining smooth transitions between video clips. Due to the high realism and dynamism of the generated clips, seamless continuity is difficult to achieve. While the method in  Fig. 8 mitigates this by using motion frames, the discrepancy between real frames used during training and synthetic frames used during inference leads to cumulative errors. As the video length increases, these errors accumulate, resulting in a degradation of quality in the later clips. Future work will focus on minimizing this error propagation, possibly through improved temporal modeling or alternative approaches to frame synthesis during inference. (3) High-Resolution: Although Latent Diffusion Models (LDM) [34] leverage VAE to enhance the resolution of synthesized images, high resolution remains a significant challenge for video tasks. Recent works in portrait animation, such as [41, 45, 47], have circumvented this issue through a two-stage training process. However, we hope that future research will explore more effective solutions for generating high-resolution outputs in end-to-end models like LetsTalk.

### v2

@@
-LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis
-Abstract
-* Equal contribution. Correspondence to Ruibo Fu: ruibo.fu@nlpr.ia.ac.cn.
-Portrait image animation using audio has rapidly advanced, enabling the creation of increasingly realistic and expressive animated faces.
-The challenges of this multimodality-guided video generation task involve fusing various modalities while ensuring consistency in timing and portrait.
-We further seek to produce vivid talking heads.
-To address these challenges, we present LetsTalk (LatEnt Diffusion TranSformer for Talking Video Synthesis), a diffusion transformer that incorporates modular temporal and spatial attention mechanisms to merge multimodality and enhance spatial-temporal consistency.
-To handle multimodal conditions, we first summarize three fusion schemes, ranging from shallow to deep fusion compactness, and thoroughly explore their impact and applicability.
-Then we propose a suitable solution according to the modality differences of image, audio, and video generation.
-For portrait, we utilize a deep fusion scheme (Symbiotic Fusion) to ensure portrait consistency.
-For audio, we implement a shallow fusion scheme (Direct Fusion) to achieve audio-animation alignment while preserving diversity.
-Our extensive experiments demonstrate that our approach generates temporally coherent and realistic videos with enhanced diversity and liveliness.
-The project page is available at Let’s Talk.
-{strip}
-[Uncaptioned image][Uncaptioned image]
-Figure 1: We propose LetsTalk, a diffusion-based transformer for audio-driven portrait image animation.
-Left: Given a single reference image and audio, LetsTalks can produce a realistic and vivid video aligned with the input audio.
-Note that each column corresponds to the same audio.
-The results show that LetsTalk can drive consistent and reasonable mouth motions for the input audio.
-Right: Generation quality vs.
-inference time on the HDTF dataset, the circle area reflects the number of parameters of the method.
-Our LetsTalk achieves the best quality while also being highly efficient in inference, compared to current mainstream diffusion-based methods such as Hallo and AniPortrait.
-In addition, our base version (LetsTalk-B) achieves performance similar to Hallo with only 8 $\times$ fewer parameters.
-1 Introduction
+Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance
+Abstract.
+Portrait image animation using audio has rapidly advanced, but challenges remain in efficiently fusing multimodal inputs while ensuring temporal and portrait consistency with minimal computational cost.
+To address this, we present LetsTalk, a LinEar diffusion TranSformer for Talking video synthesis.
+LetsTalk incorporates a deep compression autoencoder to obtain efficient latent representations, and a spatio-temporal-aware transformer with efficient linear attention to effectively fuse multimodal information and enhance spatio-temporal consistency.
+We systematically explore and summarize three fusion schemes, ranging from shallow to deep fusion.
+We thoroughly analyze their characteristics, applicability, and trade-offs, thereby bridging critical gaps in multimodal conditional guidance.
+Based on modality differences of image, audio, and video generation, we adopt deep (Symbiotic Fusion) for portrait to ensure consistency, and shallow (Direct Fusion) for audio to align animation with speech while preserving motion diversity.
+To maintain temporal consistency in long-duration video generation, we propose a memory bank mechanism that preserves inter-clip dependencies, effectively preventing degradation across extended sequences.
+Furthermore, we develop a noise-regularized training strategy that explicitly compensates for DDPM sampling artifacts, significantly improving the model’s robustness in continuous generation scenarios.
+Our extensive experiments demonstrate that our approach achieves state-of-the-art generation quality, producing temporally coherent and realistic videos with enhanced diversity and liveliness, while maintaining remarkable efficiency through its optimized model design with 8$\times$ fewer parameters.
+Talking Head Generation, Multimodal Guidance, Long-duration Generation, Linear Diffusion Transformer, Deep Compression Autoencoder
+ccs: Computing methodologies Procedural animationccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
+Figure 1.
+We introduce LetsTalk, a diffusion-based transformer for audio-driven portrait animation.
+Given a reference image and audio, LetsTalk generates realistic videos with synchronized mouth motions.
+As shown in the Left figure, each column corresponds to the same audio, demonstrating consistent and accurate lip movements.
+The Right figure compares generation quality and inference time on the HDTF dataset, where circle size represents model parameters.
+LetsTalk achieves superior quality and efficiency, outperforming methods like Hallo and AniPortrait.
+Notably, our base version (LetsTalk-B) matches Hallo’s performance with 8$\times$ fewer parameters.
+[…]
+Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
+1. Introduction
 Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio.
-Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc.
-However, producing reasonable animation faces the following main challenges: consistency in timing and portrait, audio-animation alignment, and diversity.
-To overcome these challenges, several methods [52, 45] adopt a 3D morphable model (3DMM) [4, 26, 3] to represent the reference portrait and use audio signals to drive the 3DMM.
-They feed the 3D signals (e.g.
-coefficients, poses) from the driven 3DMM into a neural renderer to generate a talking head.
-While achieving great consistency in timing and portrait, they tend to produce dull results, such as unblinking eyes and fixed hair, which damage the diversity of animations.
-In contrast to these 3D-aware methods, some other methods [7, 32, 55, 10, 54] mainly focus on capturing the correspondence between the given audio and related 2D movements, like facial muscle and lip motions, directly aligning animation with audio.
-They implicitly learn the correspondence between audio and animation in the latent space, animating the portrait image by feeding forward the audio.
-However, these methods still struggle to disentangle the components representing the static and dynamic aspects of the face, even hindering the temporal consistency and diversity of videos.
-Recently, diffusion models [18, 13, 34] have emerged as a promising technique to generate realistic 2D images [34, 51] and have even achieved remarkable results in the video generation [29, 28].
-Notably, AnimateAnyone [21] incorporates motion modules and ReferenceNet into diffusion models to generate realistic and pose-controllable human animation, which verifies that the end-to-end diffusion models can seamlessly capture the correspondence between the condition signal and human animation.
-Inspired by it, recent methods [47, 9, 11] replace the pose signals in AnimateAnyone with audio signals and achieve talking head generation.
-Leveraging diffusion models, these methods excel at generating various unconscious actions, thereby enhancing diversity.
-While achieving promising results, these methods adopt U-Net as the backbone and lack the flexibility for better fusing multimodality.
-In addition, although the U-Net backbone works well for image generation, it introduces considerable payloads and struggles with scalability for video generation.
-Therefore, it is necessary to propose a transformer framework to solve this task and reasonably fuse multimodality, ensuring better portrait consistency and producing more vivid characters.
-In this work, we present LetsTalk, a latent diffusion transformer to solve this video generation-like task via the diffusion paradigm.
-In LetsTalk, we first employ a variational autoencoder (VAE) to reduce this task into a generation task within low-resolution latent space.
-To enhance the spatial-temporal consistency, we equip each transformer block with temporal and spatial attention modules to capture intra-frame and cross-frame coherence.
-Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference image and audio sequences.
-To better exploit multimodality to guide video generation, we summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, which uses a fixed modality to guide all features in the backbone through a cross-attention mechanism; 2) Siamese Fusion, like ReferenceNet, which feeds modality into a backbone-like network and uses the corresponding features to guide the features in the backbone; 3) Symbiotic Fusion, which concatenates modality with the input at the beginning and achieves guidance only through backbone’s inherent self-attention mechanisms.
-After summarizing these fusion schemes, we explore the impact of adopting different schemes for the given reference image and audio sequence.
-To better guide video generation, we use Symbiotic Fusion for the reference image whose modality is close to the video while strongly ensuring the portrait consistency of the generated video.
-Given the audio, we ask the animation to be aligned with it, generating plausible animation, especially lip motions, and further pursuing diversity (e.g.
-blinking, head moving, and expressions).
-Considering the audio modality differs from video, we adopt the Direct Fusion that ensures the audio-animation alignments and avoids ambiguity caused by significant modality differences in the fusion.
-Additionally, we find that unconscious portrait actions, closely related to diversity, are primarily generated during the denoising of diffusion models, and the shallow fusion scheme effectively preserves the diversity.
-The empirical results indicate that our innovative framework realizes realistic and high-fidelity talking-head synthesis, advancing the frontiers of realism and expressiveness in audio-driven video generation.
-To the best of our knowledge, our method is the first work to introduce the diffusion transformer for audio-driven portrait animation.
+Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc..
+Previous works (Zhang et al., 2023a; Wei et al., 2024) use a 3D morphable model (3DMM) (Blanz and Vetter, 1999; Li et al., 2017; Bao et al., 2021) to represent reference portrait and drive them with audio signals.
+These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair.
+In contrast, audio-to-2D facial motion methods (Chen et al., 2019; Prajwal et al., 2020; Zhou et al., 2020; Cheng et al., 2022; Zhou et al., 2021) directly align audio with facial dynamics via latent space learning.
+However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
+Recently, diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022) have emerged as a powerful paradigm for generating realistic 2D images (Rombach et al., 2022; Zhang et al., 2023b) and have demonstrated remarkable success in video generation (Ma et al., 2024; Lu et al., 2023).
+Notably, AnimateAnyone (Hu, 2024) introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion.
+Building on this, recent methods (Xu et al., 2024b; Chen et al., 2024; Cui et al., 2024) have extended this approach by replacing pose signals with audio signals for talking head generation.
+These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
+Despite these advancements, several critical challenges remain unaddressed.
+First, current approaches suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific requirements; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity.
+Second, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
+Lastly, while existing methods attempt to achieve long-duration portrait animation through sequential frame synthesis, they fail to address the progressive accumulation of residual noise in generated frames.
+This phenomenon inevitably introduces cumulative artifacts (e.g., facial distortions or flickering) that amplify over extended sequences, ultimately degrading output fidelity.
+In this work, we introduce LetsTalk, a linear diffusion transformer for efficient generation of realistic, long-duration talking videos under multimodal conditions.
+Built upon a diffusion transformer as the video generation backbone, LetsTalk replaces conventional text-to-image architectures with an efficient design tailored for audio-driven animation.
+Key to our approach is a deep compression autoencoder that compresses input frames 32×, drastically reducing the number of latent tokens while preserving critical visual details.
+Each transformer block incorporates dedicated temporal and spatial modules, where linear attention (Xie et al., 2024a) ensures efficient modeling of intra-frame and cross-frame coherence without sacrificing computational efficiency.
+Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio.
+To better exploit multimodality to guide video generation, we first summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion.
+We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities, and further investigate the underlying factors influencing such adaptability—thereby bridging theoretical gaps in conditional guidance modeling.
+To better guide video generation effectively, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency due to its visual similarity to the video.
+For audio, we employ Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion).
+To enhance long-duration video generation, we incorporate a memory bank mechanism to maintain contextual continuity across sequential video sequences.
+Furthermore, we propose a novel noise-regularized training strategy and a patch drop augmentation to align the memory bank between training and inference.
+By explicitly simulating the residual noise accumulation that occurs during inference (where errors propagate as the video extends), our approach learns to compensate for such artifacts during training for robust long-sequence generation.
 We summarize the contributions of this work as follows:
-  • We present LetsTalk that pioneers the utilization of transformers in diffusion-based, audio-driven portrait animation by introducing modular spatial-temporal attention mechanisms;
-  • We summarize and deeply analyze various multimodal fusion schemes, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality;
-  • Our experiments on standard benchmarks demonstrate the superiority of our methods to other approaches in driving portrait animation via audio.
-2 Related Work
-2.1 Diffusion Models for Video Generation
-Diffusion models have demonstrated exceptional potential across various domains, particularly in image [18, 13] and video generation [17, 21, 20, 57].
-A notable example is Stable Diffusion [34], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [37].
-This approach has been extended to more complex tasks like video generation [17, 21], where models such as Video Diffusion Models (VDM)[20] apply the diffusion process with space-time factorized training, and ImagenVideo[19] enhances VDM with cascaded diffusion models for high-resolution outputs.
-Further advances, including VideoComposer [43], AnimateDiff [17], and VideoCrafter [6], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents.
-A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [31], which facilitates scaling to larger datasets and model parameters.
-Recent works, such as Latte [29] and VDT [28], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch.
+  • We present LetsTalk, a linear diffusion transformer for audio-driven portrait animation, which employs efficient linear attention for spatiotemporal modeling and integrates a deep compression autoencoder to achieve 32 $\times$ latent space reduction while preserving visual fidelity;
+  • We systematically analyze three multimodal fusion schemes, characterize their adaptation properties, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality guidance;
+  • We enable long-duration video generation via a memory bank framework with noise-regularized training, aligning memory states between training and inference to mitigate cumulative artifacts in extended sequences;
+  • Our experiments demonstrate our method’s superiority in generating realistic, temporally coherent videos with enhanced diversity, achieving an optimal balance between computational efficiency and output quality.
+2. Related Work
+2.1. Diffusion Models for Video Generation
+Diffusion models have demonstrated exceptional potential across various domains, particularly in image (Ho et al., 2020; Dhariwal and Nichol, 2021) and video generation (Guo et al., 2023; Hu, 2024; Ho et al., 2022b; Zhu et al., 2024).
+A notable example is Stable Diffusion (Rombach et al., 2022), which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets (Schuhmann et al., 2022).
+This approach has been extended to more complex tasks like video generation (Guo et al., 2023; Hu, 2024), where models such as Video Diffusion Models (VDM)(Ho et al., 2022b) apply the diffusion process with space-time factorized training, and ImagenVideo(Ho et al., 2022a) enhances VDM with cascaded diffusion models for high-resolution outputs.
+Further advances, including VideoComposer (Wang et al., 2024b), AnimateDiff (Guo et al., 2023), and VideoCrafter (Chen et al., 2023b), focus on synthesizing videos by conditioning the diffusion process on images and blending image latents.
+A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), which facilitates scaling to larger datasets and model parameters.
+Recent works, such as Latte (Ma et al., 2024) and VDT (Lu et al., 2023), have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch.
 Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
-2.2 Audio-driven Portrait Image Animation
+2.2. Audio-driven Portrait Image Animation
 In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs.
-Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [32] achieving high accuracy for specific identities.
-Subsequently, methods such as ATVGnet [7] and MakeItTalk [55] introduced facial landmarks and disentangled representations to improve video frame generation.
-The integration of 3D information in models like SadTalker [52], Genface [50], and DiffTalk [38] further improved head movements and expressions.
-A significant advancement came with the introduction of diffusion models, including Diffused Heads [39], DreamTalk [30], and VividTalk [40], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles.
-Recent innovations like VASA-1 [48], EMO [41] and AniPortrait [45] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements.
-AniTalker [27] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars.
-LivePortrait [16] extends the implicit-keypoint framework, balancing efficiency and controllability.
-Most recently, Hallo [47, 11] and X-Portrait [46] have adopted the latent diffusion framework pioneered in AnimateAnyone [21] or ControlNet [51], adding more control over elements such as lip movements, expressions, and poses.
-2.3 Conditionally Guided Diffusion Model
+Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert (Prajwal et al., 2020) achieving high accuracy for specific identities.
+Subsequently, methods such as ATVGnet (Chen et al., 2019) and MakeItTalk (Zhou et al., 2020) introduced facial landmarks and disentangled representations to improve video frame generation.
+The integration of 3D information in models like SadTalker (Zhang et al., 2023a), Genface (Ye et al., 2023), and DiffTalk (Shen et al., 2023) further improved head movements and expressions.
+A significant advancement came with the introduction of diffusion models, including Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), and VividTalk (Sun et al., 2023), which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles.
+Recent innovations like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2024) and AniPortrait (Wei et al., 2024) have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements.
+AniTalker (Liu et al., 2024) introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars.
+LivePortrait (Guo et al., 2024) extends the implicit-keypoint framework, balancing efficiency and controllability.
+Most recently, Hallo (Xu et al., 2024b; Cui et al., 2024), EchoMimic (Chen et al., 2024) and X-Portrait (Xie et al., 2024b) have adopted the latent diffusion framework pioneered in AnimateAnyone (Hu, 2024), adding more control over elements such as lip movements, expressions, and poses.
+2.3. Conditionally Guided Diffusion Model
 The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research.
-Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [33, 1, 5, 49, 24].
-ControlNet [51] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning.
-In video generation, AnimateAnyone [21] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images.
-Gen-1 [15] integrates structural guidance through depth maps and cross-attention, while MCDiff [8], LaMD [22], and MotionCtrl [44] focus on motion-guided video generation, enhancing the realism of human actions.
-For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [38, 39].
-Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [41, 47].
-While control over both images and videos in latent diffusion models has seen significant advancements, diffusion transformers remain underdeveloped in this area, particularly when it comes to handling multiple modalities (e.g.
-image, video, audio).
-To address this gap, we explore various diffusion conditioning strategies tailored to the transformer architecture, aiming to achieve better control and enhanced generation quality.
-(a) Overview of LetsTalk
-(b) Transformer Block
-Figure 2: The overview of our method (a) and the illustration of our designed transformer block (b).
-For better illustration, we omit the timestep encoder and Layer Norm in (b).
-LetsTalk integrates transformer blocks equipped with both temporal and spatial attention modules, designed to capture intra-frame spatial details and establish temporal correspondence across time steps.
+Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms (Ramesh et al., 2022; Avrahami et al., 2022; Brooks et al., 2023; Yang et al., 2024; Khachatryan et al., 2023).
+ControlNet (Zhang et al., 2023b) enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning.
+In video generation, AnimateAnyone (Hu, 2024) introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images.
+Gen-1 (Esser et al., 2023) integrates structural guidance through depth maps and cross-attention, while MCDiff (Chen et al., 2023a), LaMD (Hu et al., 2023), and MotionCtrl (Wang et al., 2024a) focus on motion-guided video generation, enhancing the realism of human actions.
+For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits (Shen et al., 2023; Stypułkowski et al., 2024).
+Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information (Tian et al., 2024; Xu et al., 2024b).
+While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio).
+To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
+Figure 2.
+The illustration of our method.
+LetsTalk integrates a deep compression autoencoder that efficiently reduces spatial redundancy while preserving temporal features, and transformer blocks equipped with dual temporal and spatial attention modules, enabling the capture of intra-frame spatial details while establishing temporal correspondences across frames.
 After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding.
 Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding.
-3 Method
-We present LetsTalk, a diffusion transformer for portrait image animation driven by speech, as shown in Fig.
+To support long-duration video generation, we introduce a memory bank module to preserve temporal consistency across video sequences.
+Additionally, we devise a novel training strategy to to align memory bank between training and inference.
+By explicitly simulating the residual noise during inference, our approach learns to compensate for such artifacts, ensuring robust long-duration generation.
+3. Method
+We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig.
 2.
-In this section, we first briefly introduce diffusion models and the formulation of our task in Sec.
+In this section, we first detail our efficient transformer architecture in Sec.
 3.1.
-Then, we delve into the details of our transformer backbone in Sec.
-3.2.
-Finally, we summarize multimodal fusion schemes and present our adopted strategy for fusing the input reference portrait and audio in Sec.
+Next, we outline multimodal fusion schemes in Sec.
+3.2 and describe our strategy for integrating reference portraits with audio in Sec.
 3.3.
-3.1 Preliminaries
-Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters.
-By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,1)$, forming a Markov chain over $T$ steps.
-Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$.
-This is optimized via the variational lower bound of $\log p(x_{0})$:
-$\begin{split}\vspace{-0.3cm}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
-With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances.
-Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
-$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
-To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
-Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework.
-DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches.
-This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently.
-Different from ViT, DiT incorporates a variational autoencoder (VAE) [25] in the framework, thus runs in the latent space, rather than pixel space.
-Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
-Since we follow DiT and introduce a pretrained VAE in our framework, our goal is reduced from generating a video $\bm{V}\in\mathbb{R}^{F\times\hat{H}\times\hat{W}\times 3}$ to generating a latent feature $\bm{F}\in\mathbb{R}^{F\times H\times W\times C}$, where $F$ denotes the number of video frames333 The encoder in VAE downsamples the image by a factor of $8$, and $\hat{H}=8\times H,\hat{W}=8\times W$ We first separate the latent feature $\bm{F}$ into non-overlapping square patches along the spatial dimensions (i.e., $H$ and $W$) and obtain a patch sequence of length $F\times P$.
-For each patch, we flatten it and use an MLP to project it into the latent space as $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, where $d$ is the dimension of the latent space.
-Then we add spatial and temporal positional embeddings [42] to the latent feature and make it sensitive to spatial and temporal information.
-Finally, we feed the processed latent feature into transformers and perform denoising.
-(a) Direct Fusion
-(b) Siamese Fusion
-(c) Symbiotic Fusion
-Figure 3: Illustration of three multimodal fusion schemes, our transformer backbone is formed by the left-side blocks.
-(a) Direct Fusion.
-Directly feeding condition into each block’s cross-attention module; (b) Siamese Fusion.
-Maintaining a similar transformer and feeding the condition into it, extracting the corresponding features to guide the features in the backbone; (c) Symbiotic Fusion.
-Concatenating modality with the input at the beginning, then feeding it into the backbone, achieving fusion via the inherent self-attention mechanisms.
-3.2 Efficient Spatial-Temporal-Aware Transformer Backbone
-Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we feed it into a transformer and predict the noise to be removed.
-For video generation, we extend the backbone in DiT that only consists of spatial transformer blocks.
-It is straightforward to flatten the $\bm{Z}$ into $F\times P$ tokens with shape $(F\times P,d)$ and repeatedly conduct a self-attention mechanism to achieve spatiotemporal fusion.
-However, this approach results in a huge payload for computing the attention matrix with shape $(F\times P,F\times P)$, as it grows with the square of the increase in resolution or frames.
-To address this problem, we propose to achieve spatiotemporal fusion in an “interleaved fusion” manner, implemented by alternately stacking spatial and temporal blocks.
-Considering that $F$ refers to the temporal dimension (i.e.
-the number of frames) and $P$ refers to the spatial dimension (i.e.
-the number of patches in each frame), to feed the feature into the spatial block, we maintain the feature with shape $(F,P,d)$ and calculate the attention matrix with shape $(F,P,P)$ to capture spatial information exclusively among patches in the same frame.
-Similarly, to feed the feature into the temporal block, we maintain the feature with shape $(P,F,d)$ and calculate the attention matrix with shape $(P,F,F)$ to capture temporal information among frames in the same patch.
-Compared with the straightforward scheme that produces an attention matrix with $F^{2}\times P^{2}$ parameters, our scheme significantly reduces the amount to $F\times P^{2}+P\times F^{2}$.
-Our experiments verify the efficacy of our backbone extension in terms of image quality and temporal consistency.
-3.3 Multimodal Fusion Schemes for Portrait and Speech Audio
+Finally, we introduce our noise-regularized training approach for long-duration video generation Sec.
+3.4.
+3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
+Deep Compression Autoencoder: Building upon latent diffusion models, we employ a deep compression autoencoder (Xie et al., 2024a) that downsamples images by a factor of 32 (vs.
+8 in conventional VAEs), while maintaining reconstruction quality comparable to SDXL’s autoencoder through advanced architectural designs.
+This design is motivated by two key insights: (1) high-resolution video frames contain substantial redundant information that can be safely compressed, and (2) for audio-driven generation, separating compression from denoising enables specialized processing - the autoencoder preserves crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
+The compressed latent feature $\mathcal{F}\in\mathbb{R}^{F\times H\times W\times C}$ is processed through non-overlapping patches, where spatial-temporal positional embeddings (Vaswani, 2017) preserve structural relationships.
+This compact representation enables efficient transformer processing while reducing memory requirements by 4$\times$ compared to conventional approaches.
+Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal.
+While the original DiT (Peebles and Xie, 2023) processes individual frames independently, our video generation framework extends this through interleaved spatiotemporal fusion.
+The naive approach of flattening $\bm{Z}$ into $(F\times P,d)$ tokens with standard attention incurs prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution video generation.
+Our architecture employs a dual-block decomposition strategy that alternates between spatial and temporal modeling, significantly reducing computational complexity from $\mathcal{O}(F^{2}P^{2})$ to $\mathcal{O}(F^{2}+P^{2})$.
+To reduce computational complexity, we replace the traditional softmax attention with ReLU linear attention (Katharopoulos et al., 2020).
+For spatial modeling, the input is first reshaped into $(F,P,d)$ dimensions to compute intra-frame attention using:
+(1) $\bm{Z_{s}}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}$
+where $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$, $\phi(x)=\text{ReLU}(x)$.
+The linear attention mechanism further optimizes this by reusing shared terms across all queries: $\sum_{j=1}^{N}\text{ReLU}(\bm{K}_{j})\bm{V}_{j}^{T}\in\mathbb{R}^{d\times d}$.
+For temporal modeling, the tensor to $(P,F,d)$ dimensions.
+Since $F\ll P$, we preserve the expressive power of standard softmax attention for inter-frame interactions: $\bm{Z_{t}}=\text{Attn}(\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t})=\frac{\bm{V}_{t}(\bm{K}_{t}^{\top}\bm{Q}_{t})}{\sqrt{d}}$.
+Through this hybrid decomposition with linear attention, the overall complexity is reduced to $\mathcal{O}(F^{2}+Pd))$ achieving near-linear efficiency while preserving global receptive fields.
+Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance.
+Following SANA (Xie et al., 2024a), we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU.
+The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality.
+Our experiments demonstrate this architecture achieves superior computational efficiency without compromising visual quality or temporal consistency compared to quadratic-complexity baselines in Fig.
+1.
+Figure 3.
+Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention.
+The backbone architecture (left-side blocks) remains consistent across all approaches.
+Schemes Integration State Params Modality Adapt.
+Direct Cross-attention Static Dedicated Weak alignment
+Siamese Self-attention/Add Dynamic Dedicated Flexible
+Symbiotic Self-attention Dynamic Shared Strong
+Table 1. Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
+3.2. Analysis of Multimodal Fusion Schemes
 After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task.
 Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
-  • Direct Fusion (Fig. 3(a)) directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention.
-  • Siamese Fusion (Fig.
-3(b)) feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
-  • Symbiotic Fusion (Fig. 3(c)) concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
-After defining these schemes with fusion compactness ranging from shallow to deep, we aim to adopt reasonable solutions for input modalities (i.e. the reference image of portrait and audio).
-Given the reference image, we require that the character of the generated video is consistent with the reference, i.e.
-using the reference image to strongly guide video generation.
-Moreover, Since the image modality is close to the video, Symbiotic Fusion is appropriate.
+  • Direct Fusion directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention;
+  • Siamese Fusion feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
+  • Symbiotic Fusion concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
+Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization.
+Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models (Rombach et al., 2022; Nichol et al., 2021).
+In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data (Wang et al., 2024c).
+Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
+Condition States: The static versus dynamic nature of conditions reflects their constraint strength.
+Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet (Hu, 2024) and ControlNet (Zhang et al., 2023b) that multi-level conditioning enhances generation controllability.
+This flexibility improves feature alignment and model performance.
+Parameters Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters.
+Direct Attention, relying solely on cross-attention, also does not share parameters.
+Modality Adaptation: The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks.
+Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios.
+Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment.
+This is exemplified in models such as ControlNet (Zhang et al., 2023b), which employs visual cue-level alignment, and AnimateAnyone (Hu, 2024), which utilizes pixel-level alignment.
+The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand.
+Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters.
+This makes it particularly effective for large-scale applications, such as AnimateAnyone (Hu, 2024), where shared parameters ensure robust and efficient multimodal integration.
+The comparative analysis of the three fusion approaches are summarized in Tab.
+1
+3.3. Guidance on Reference Portrait and Audio
+Given the reference image, we require that the character of the generated video is consistent with the reference.
+Since the image modality is close to the video, Symbiotic Fusion proves appropriate.
 Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space.
-Then we temporally repeat the portrait feature $\bm{P}$ with $F$ (i.e.
-frames) times and concatenate it with the latent feature $\bm{F}$ along the spatial dimension [^4] To concatenate latent feature $\bm{F}\in\mathbb{R}^{F\times P\times d}$ with the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$, we extend and repeat the portrait feature along the temporal dimension as $\tilde{\bm{P}}\in\mathbb{R}^{F\times P\times d}$, the shape of concatenated latent feature is $[F,2P,d]$.
-and feed them into the transformer backbone.
-In contrast, we observe that Direct Fusion attempts to use shallow portrait feature $\bm{P}$ to guide the learning of all blocks, leading to misaligning the features in the deeper blocks.
-While Siamese Fusion addresses this misalignment and achieves comparable results to Symbiotic Fusion, it needs to maintain a siamese transformer similar to the backbone, which introduces additional model parameters and complicates the learning task.
-Datasets Raw Filtered
-#IDs #Hours #IDs #Hours
-HDTF 411 15.75 356 13.95
-CelebV-HQ 35,666 68 22,423 41.27
-Train - - 20,501 49.69
-Eval - - 2,278 5.522
-Figure 4: The dataset splits for HDTF and CelebV-HQ, including raw/filtered data and training/evaluation distributions.
-As for the audio, we require it to reasonably drive the animation, especially the lip motion [52].
+Then we temporally repeat the portrait feature $\bm{P}$ with $F$ frames and concatenate it with the latent feature $\bm{Z_{s}}\in\mathbb{R}^{F\times P\times d}$ along the spatial dimension, which can be formulated as: $\bm{Z}=\text{concat}\left(\,\bm{Z_{s}},\,\tilde{\bm{P}}\,\right)$ denotes the temporally extended portrait feature.
+This operation yields a concatenated latent feature $\bm{Z}\in\mathbb{R}^{F\times 2P\times d}$.
+By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames.
+In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers.
+While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
+As for the audio, we require it to reasonably drive the animation, especially the lip motion (Zhang et al., 2023a).
 However, certain unconscious actions (e.g.
 blinking and head movements) are almost independent of the audio yet serve to diversity.
-Therefore, we aim for the audio to have a minimal correlation with these animations.
 Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity.
 In a nutshell, Direct Fusion is concise yet effective for fusing the input audio.
-In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [47] and extract richer semantic information from different layers of the wav2vec encoder [36].
-A sliding window mechanism is employed to capture long-term sequential audio information.
-Through nonlinear mapping, we obtain audio tokens enriched with semantic content, which are aligned with latent features in the same feature space.
+In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow (Xu et al., 2024b) and extract richer semantic information from different layers of the wav2vec encoder (Schneider et al., 2019).
+Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space.
 Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation.
@@
 4.4 verify the efficacy of our mechanism in feature fusion.
-4 Experiment
-4.1 Experimental Setups
-Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [53] and CelebV-HQ [56].
+Figure 4. The illustration of the long-duration generation.
+3.4. Long-Duration Generation
+Unlike the previous frame-by-frame generation approach (Shen et al., 2023; Stypułkowski et al., 2024), our end-to-end diffusion transformer generates a video clip in a single sampling process.
+In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences.
+Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible.
+To address this, we introduce a memory bank that stores information from previously generated clips.
+This data undergoes patchification and temporal transpose operations, transforming it into $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ represents the number of frames selected from the prior clips.
+Leveraging the scalability of the Transformer architecture, the memory bank directly provides prior information through Symbiotic Fusion with $\bm{Z_{t}}$.
+However, the inherent distribution discrepancy between training (real data) and inference (synthetic data) poses a fundamental challenge for diffusion models, with error accumulation becoming increasingly severe as video length grows.
+To mitigate this latent risk, we develop a noise-regularized training strategy.
+Specifically, we first establish that diffusion models inherently retain residual noise even after finite-step denoising due to the discrete approximation of continuous noise schedules, which progressively accumulates as the video sequence extends.
+Building on the insight, we introduce a bounded stochastic noiser that perturbs the memory bank contents with controlled Gaussian noise $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$ (see Fig.
+4) into the memory bank during training.
+Inspired by Hallo2 (Cui et al., 2024), we further apply a patch-drop data augmentation strategy to the memory bank.
+Unlike Hallo2 which protects face patches from dropping, we employ completely random dropping to enhance the model’s robustness to the memory bank.
+This deliberate corruption serves two purposes: (1) it faithfully simulates inference-time distribution shifts, and (2) it preserves crucial temporal dependencies.
+The resulting framework demonstrates significantly improved robustness against error propagation in long-duration generation tasks.
+Figure 5. The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
+4. Experiment
+4.1. Experimental Setups
+Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022).
 These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds.
-To ensure high-quality training data, we adopt the preprocess scheme used in [47].
+To ensure high-quality training data, we adopt the preprocess scheme used in (Xu et al., 2024b).
 We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots.
 Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence.
-After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 22,423 clips with 41.27 hours from the CelebV-HQ dataset.
+After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset.
 During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
 Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID).
-Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [23].
+Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs (Karras et al., 2019).
 SSIM indices measure the structural similarity between ground truth videos and generated videos.
 E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure.
-It extracts facial expression parameters via a reconstruction method [12] and computes FID to quantify the gap between generated and ground truth expressions.
-Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [30], SadTalker (138M) [52], AniPortrait (2.8B) [45], and Hallo (2.4B) [47].
-The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes, as shown in Fig.
-4.
-Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
-Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
+It extracts facial expression parameters via a reconstruction method (Deng et al., 2019) and computes FID to quantify the gap between generated and ground truth expressions.
+Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) (Ma et al., 2023), SadTalker (138M) (Zhang et al., 2023a), AniPortrait (2.8B) (Wei et al., 2024), Hallo (2.4B) (Xu et al., 2024b), Hallo (2.4B) (Cui et al., 2024) and EchoMimic (2.6B) (Chen et al., 2024).
+The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
+Figure 6. The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
 Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
@@
 AniPortrait 78.284 331.117 0.623 2.903
+EchoMimic 55.791 299.343 0.639 2.493
 Hallo 45.754 285.988 0.698 2.235
+Hallo2 44.343 241.628 0.657 2.017
 Ours (Base) 46.240 289.385 0.666 2.587
-Ours (Large) 37.134 219.119 0.632 1.962
-Table 1: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
+Ours (Large) 31.761 219.119 0.632 1.962
+Table 2. The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
+Implementation details Our approach, in contrast to previous methods (Tian et al., 2024; Xu et al., 2024b; Wei et al., 2024) that rely on a two-stage training process, adopts an end-to-end training strategy.
+We set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
+4.2. Quantitative Results
+The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab.
+2 and Tab.
+3, respectively.
+The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment.
+Notably, our approach achieves the lowest FID (31.761) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset.
+These results show that our method generates both realistic and temporally consistent talking-head videos.
+For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
+4.3. Qualitative Results
+We have visualized the results of different methods on the HDTF dataset (Fig.
+8) and the CelebV-HQ dataset (Fig.
+9).
+From the visualization results, it is evident that our method performs exceptionally well on both datasets.
+In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth.
+Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency.
+Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
 Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
@@
 AniPortrait 33.408 107.236 0.578 3.722
-Hallo 8.504 49.431 0.580 3.280
+EchoMimic 12.932 59.944 0.567 3.671
+Hallo 8.504 46.431 0.580 3.280
+Hallo2 8.734 41.213 0.543 3.128
 Ours (Base) 9.585 36.159 0.598 3.349
 Ours (Large) 7.610 33.175 0.584 3.038
-Table 2: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
-Implementation details Our approach differs from previous methods [41, 47, 45] that require a two-stage training process.
-We adopt an end-to-end training strategy.
-Thanks to the scalability of transformers, we set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
-Following common practices in generative modeling works [31, 2], an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model.
-Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training.
-Please refer to Sec.
-7 for more implementation details.
-4.2 Quantitative Results
-The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab.
-1 and Tab.
-2, respectively.
-The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment.
-Notably, our approach achieves the lowest FID (37.134) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset.
-These results show that our method generates both realistic and temporally consistent talking-head videos.
-For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
-4.3 Qualitative Results
-We have visualized the results of different methods on the HDTF dataset (Fig.
-9) and the CelebV-HQ dataset (Fig.
-10).
-From the visualization results, it is evident that our method performs exceptionally well on both datasets.
-In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth.
-Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency.
-Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
-More results can be found in Sec.
-8.
-4.4 Ablation Study
+Table 3. The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
+Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
+Direct Fusion 52.999 389.01 0.497 3.149
+Siamese Fusion 47.779 286.692 0.651 2.597
+Symbiotic Fusion 46.240 289.385 0.666 2.587
+Table 4. Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
+Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
+Symbiotic Fusion 51.880 403.628 0.471 3.398
+Siamese Fusion 48.041 318.637 0.585 2.706
+Direct Fusion 46.240 289.385 0.666 2.587
+Table 5. Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
+4.4. Ablation Study
 In this section, we analyze the impact of different fusion schemes as illustrated in Fig.
@@
 The parameters of these transformers are simultaneously updated during training.
-We conduct experiments on the HDTF dataset using the LetsTalk-B model to validate the impact of different fusion schemes.
 Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used.
 We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab.
-3.
+4.
 Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality.
@@
 Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity.
-Audio-animation alignment refers to specific visual cues [52] (e.g.
+Audio-animation alignment refers to specific visual cues (Zhang et al., 2023a) (e.g.
 facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g.
@@
 5 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
-To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^5] Filling audio embeddings with 00 for fusion.
-and compare the diversity of character actions in the generated videos.
-The results in Fig.
-7 demonstrate that the deeper the fusion compactness, the more static the generated video.
-The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
-Figure 7: Portrait animation results without audio guidance for three fusion schemes.
-Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
-Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-Direct Fusion 52.999 389.01 0.497 3.149
-Siamese Fusion 47.779 286.692 0.651 2.597
-Symbiotic Fusion 46.240 289.385 0.666 2.587
-Table 3: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
-Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-Symbiotic Fusion 51.880 403.628 0.471 3.398
-Siamese Fusion 48.041 318.637 0.585 2.706
-Direct Fusion 46.240 289.385 0.666 2.587
-Table 4: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
-5 Conclusion
-We present LetsTalk, a novel multimodality-guided diffusion transformer to achieve superior audio-driven portrait animation results using audio input.
-In LetsTalk, we introduce modular spatial-temporal attention mechanisms into our transformer backbone to enhance temporal consistency.
-We summarize three fusion schemes to handle multimodality and deeply analyze their impact.
-We then adopt the optimal fusion schemes for input reference portrait and audio, respectively.
-Our extensive experiments demonstrate the effectiveness of LetsTalk in achieving state-of-the-art portrait animation results with a good balance between performance and payload.
-Supplementary Material
-6 Network Architecture
-6.1 Backbone
-In LetsTalk, we borrow the Variational AutoEncoder (VAE) [25] from Stable Diffusion [34].
+Method FID $\downarrow$ FVD $\downarrow$
+Baseline model with Memory Bank 52.243 276.16
++ Regularized Training (w/o patch drop) 32.746 264.709
++ Regularized Training (w patch drop) 34.821 251.684
+Table 6.
+Experimental results of long video generation.
+Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
+4.5. Long-Duration Generation
+To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation.
+Our results in Tab.
+6 demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality.
+The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709).
+On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
+5. Conclusion
+We introduced LetsTalk, an efficient linear diffusion transformer for high-fidelity talking head synthesis, tackling core challenges in multimodal fusion and long-term temporal coherence.
+Our approach integrates a deep compression autoencoder for compact latent representation, alongside a systematic exploration of fusion strategies—yielding an optimal framework with Symbiotic Fusion for portrait consistency and Direct Fusion for audio-video synchronization.
+The memory bank mechanism and noise-regularized training further mitigate error accumulation in extended sequences.
+Extensive experiments demonstrate LetsTalk’s state-of-the-art performance, generating lifelike, temporally coherent animations with enhanced diversity, all while maintaining superior efficiency through its optimized architecture.
+Appendix A Preliminaries
+Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters.
+By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
+Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$.
+This is optimized via the variational lower bound of $\log p(x_{0})$:
+(2) $\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$
+With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances.
+Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
+(3) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$
+To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
+Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework.
+DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches.
+This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently.
+Different from ViT, DiT incorporates a variational autoencoder (VAE) (Kingma, 2013) in the framework, thus runs in the latent space, rather than pixel space.
+Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
+Model Patch size Layer num N Hidden size D Paramaters
+LetsTalk-B 1 12 768  [^2] M
+LetsTalk-L 1 24 1024 773
+Table 7. Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
+Appendix B Network Architecture Details
+B.1. Backbone
+In LetsTalk, we borrow the Variational AutoEncoder (VAE) (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022).
 Given the reference portrait image and noisy video, we first encode them using the encoder in VAE.
@@
 Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab.
-5.
+7.
 Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$.
@@
 LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
-In the spatial-temporal-aware block, we integrate the block details into our pipeline to model the relationship between temporal and spatial attention.
-Inspired by  [28], we introduce a fully connected layer between the temporal and spatial attention mechanisms.
-This layer applies a linear transformation to the output of the temporal attention, enhancing the learning of temporal features and ensuring consistent dimensionality for subsequent processing.
-Model Patch size Layer num N Hidden size D Paramaters
-LetsTalk-B 2 12 768 193M
-LetsTalk-L 2 24 1024 663M
-Table 5: Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
-6.2 Audio Encoder
+B.2. Adaptive Integration
+Following VDT(Lu et al., 2023), we insert an adaptive full-connected layer between attention layers:
+(4) $\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$
+where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
+B.3. Audio Encoder
 We resample the audio signal to a 16 KHz sample rate.
-For audio feature extraction, we use wav2vec [36], employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation.
+For audio feature extraction, we use wav2vec (Schneider et al., 2019), employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation.
 To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips.
@@
 Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
-6.3 Siamese Transformer
-The Siamese Fusion scheme has been adopted in AnimateAnyone [21], EMO [41], Hallo [47], and so on.
+B.4. Siamese Transformer
+The Siamese Fusion scheme has been adopted in AnimateAnyone (Hu, 2024), EMO (Tian et al., 2024), Hallo (Xu et al., 2024b), and so on.
 The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise.
 Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network.
-Considering LetsTalk as a Diffusion Transformer, we introduce a ViT [14] pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image.
+Considering LetsTalk as a Diffusion Transformer, we introduce a ViT (Dosovitskiy, 2020) pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image.
 For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
-7 Additional Implementation Details
-7.1 Diffusion
+Figure 7.
+Portrait animation results without audio guidance for three fusion schemes.
+Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
+B.5. Impact of Audio-guidance on Diversity
+To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 00 for fusion.
+and compare the diversity of character actions in the generated videos.
+The results in Fig.
+7 demonstrate that the deeper the fusion compactness, the more static the generated video.
+The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
+Appendix C Additional Implementation Details
+C.1. Diffusion
 We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE.
 After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder.
-We adopt the hyperparameters from ADM [13], specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
+We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
 Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
-7.2 Long-Duration Generation
-Figure 8: The illustration of the long-duration generation process.
-Unlike the previous frame-by-frame generation approach [38, 39], our end-to-end diffusion transformer generates a video clip in a single sampling process.
-In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences.
-Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible.
-To address this, we follow [47] and pick the last $N$ frames from the last generated clip as the initial frames of the current clip ($N=2$ in our experiments).
-The initial frames are solely involved in guiding the temporal attention, ensuring continuity between consecutive clips during sampling.
-Fig.
-8 illustrates the long-duration generation process.
-7.3 Training
-We initialize LetsTalk from a pre-trained Latte [29] model on the FaceForensics dataset [35].
+C.2. Training
+We initialize LetsTalk from a pre-trained Latte (Ma et al., 2024) model on the FaceForensics dataset (Rössler et al., 2018).
 Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion.
@@
 Each model is trained for 300k to 500k iterations.
+Following common practices in generative modeling works (Peebles and Xie, 2023; Bao et al., 2023), an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model.
+Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training.
 Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
-7.4 Inference
+C.3. Inference
 During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio.
@@
 We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
-8 Qualitative Comparisons
+C.4. Diffusion
+We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE.
+After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder.
+We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
+Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
+Appendix D Qualitative Comparisons
 In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk.
-We show more results on the HDTF [53] and Celeb-HQ [56] benchmarks ( Fig.
+We show more results on the HDTF (Zhang et al., 2021) and Celeb-HQ (Zhu et al., 2022) benchmarks ( Fig.
 8 and Fig.
-10).
+9).
 We show the portrait animation results of LetsTalk driven by English (Fig.
-11), Chinese (Fig.
-12), and singing audios (Fig.
+10), Chinese (Fig.
+11), and singing audios (Fig.
+12).
+We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig.
 13).
-We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig.
-14).
 Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
-Figure 9: The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
-Figure 10: The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
-9 Visualization
+Figure 8. The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
+Figure 9. The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
+Appendix E Visualization
 We provide video visualizations in our attachment.
-Notably, the model used to generate the video is only trained on HDTF [53] and CelebV-HQ [56], which only include English conversation scenarios of real portraits.
+Notably, the model used to generate the video is only trained on HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022), which only include English conversation scenarios of real portraits.
 In addition, we use Chinese audio and singing audio to drive real portraits.
@@
 The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
-Figure 11: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 12: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 13: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 14: Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.
-10 Limitation and Future Work
-Despite the advancements proposed in this study regarding portrait image animation, several limitations necessitate further exploration and consideration.
-(1) Resource and Dataset Limitations: Unlike recent works such as Hallo [47] and EMO [41], which use diverse real-world datasets, our study is limited by the availability of publicly accessible datasets.
-These datasets are relatively small and may not be as robust, thus limiting the generalizability of LetsTalk.
-Additionally, the computational resources available were insufficient to fully exploit larger models, restricting us to smaller-scale models with fewer parameters.
-Future research could benefit from larger datasets and scaling up model sizes to explore the scalability of LetsTalk.
-(2) Long-Duration Animation: A key limitation in generating long-duration animations with our end-to-end framework is the challenge of maintaining smooth transitions between video clips.
-Due to the high realism and dynamism of the generated clips, seamless continuity is difficult to achieve.
-While the method in  Fig.
-8 mitigates this by using motion frames, the discrepancy between real frames used during training and synthetic frames used during inference leads to cumulative errors.
-As the video length increases, these errors accumulate, resulting in a degradation of quality in the later clips.
-Future work will focus on minimizing this error propagation, possibly through improved temporal modeling or alternative approaches to frame synthesis during inference.
-(3) High-Resolution: Although Latent Diffusion Models (LDM) [34] leverage VAE to enhance the resolution of synthesized images, high resolution remains a significant challenge for video tasks.
-Recent works in portrait animation, such as [41, 45, 47], have circumvented this issue through a two-stage training process.
-However, we hope that future research will explore more effective solutions for generating high-resolution outputs in end-to-end models like LetsTalk.
+Figure 10. English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
+Figure 11. Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
+Figure 12. Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
+Figure 13. Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.

### v3

@@
 Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance
-Abstract.
-Portrait image animation using audio has rapidly advanced, but challenges remain in efficiently fusing multimodal inputs while ensuring temporal and portrait consistency with minimal computational cost.
-To address this, we present LetsTalk, a LinEar diffusion TranSformer for Talking video synthesis.
-LetsTalk incorporates a deep compression autoencoder to obtain efficient latent representations, and a spatio-temporal-aware transformer with efficient linear attention to effectively fuse multimodal information and enhance spatio-temporal consistency.
-We systematically explore and summarize three fusion schemes, ranging from shallow to deep fusion.
-We thoroughly analyze their characteristics, applicability, and trade-offs, thereby bridging critical gaps in multimodal conditional guidance.
-Based on modality differences of image, audio, and video generation, we adopt deep (Symbiotic Fusion) for portrait to ensure consistency, and shallow (Direct Fusion) for audio to align animation with speech while preserving motion diversity.
-To maintain temporal consistency in long-duration video generation, we propose a memory bank mechanism that preserves inter-clip dependencies, effectively preventing degradation across extended sequences.
-Furthermore, we develop a noise-regularized training strategy that explicitly compensates for DDPM sampling artifacts, significantly improving the model’s robustness in continuous generation scenarios.
-Our extensive experiments demonstrate that our approach achieves state-of-the-art generation quality, producing temporally coherent and realistic videos with enhanced diversity and liveliness, while maintaining remarkable efficiency through its optimized model design with 8$\times$ fewer parameters.
-Talking Head Generation, Multimodal Guidance, Long-duration Generation, Linear Diffusion Transformer, Deep Compression Autoencoder
-ccs: Computing methodologies Procedural animationccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
-Figure 1.
-We introduce LetsTalk, a diffusion-based transformer for audio-driven portrait animation.
-Given a reference image and audio, LetsTalk generates realistic videos with synchronized mouth motions.
-As shown in the Left figure, each column corresponds to the same audio, demonstrating consistent and accurate lip movements.
-The Right figure compares generation quality and inference time on the HDTF dataset, where circle size represents model parameters.
-LetsTalk achieves superior quality and efficiency, outperforming methods like Hallo and AniPortrait.
-Notably, our base version (LetsTalk-B) matches Hallo’s performance with 8$\times$ fewer parameters.
+Abstract
 […]
-Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
-1. Introduction
+Index Terms:
+Talking Head Generation, Long-duration Generation, Multimodal Guidance, Linear Diffusion Transformer, Deep Compression Autoencoder
+[…]
+I Introduction
 Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio.
-Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc..
-Previous works (Zhang et al., 2023a; Wei et al., 2024) use a 3D morphable model (3DMM) (Blanz and Vetter, 1999; Li et al., 2017; Bao et al., 2021) to represent reference portrait and drive them with audio signals.
+Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc.
+Despite recent progress, long-duration talking video synthesis still faces persistent challenges in simultaneously achieving high video quality, portrait and temporal consistency, and computational efficiency.
+As video length increases, issues such as visual degradation, loss of identity consistency, temporal incoherence, and error accumulation become increasingly prominent, severely impacting the realism and reliability of generated results.
+Previous works [1, 2] use a 3D morphable model (3DMM) [3, 4, 5] to represent reference portraits and drive them with audio signals.
 These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair.
-In contrast, audio-to-2D facial motion methods (Chen et al., 2019; Prajwal et al., 2020; Zhou et al., 2020; Cheng et al., 2022; Zhou et al., 2021) directly align audio with facial dynamics via latent space learning.
+In contrast, audio-to-2D facial motion methods [6, 7, 8, 9, 10] directly align audio with facial dynamics via latent space learning.
 However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
-Recently, diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022) have emerged as a powerful paradigm for generating realistic 2D images (Rombach et al., 2022; Zhang et al., 2023b) and have demonstrated remarkable success in video generation (Ma et al., 2024; Lu et al., 2023).
-Notably, AnimateAnyone (Hu, 2024) introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion.
-Building on this, recent methods (Xu et al., 2024b; Chen et al., 2024; Cui et al., 2024) have extended this approach by replacing pose signals with audio signals for talking head generation.
+Recently, diffusion models [11, 12, 13] have emerged as a powerful paradigm for generating realistic 2D images [13, 14] and have demonstrated remarkable success in video generation [15, 16].
+Notably, AnimateAnyone [17] introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion.
+Building on this, recent methods [18, 19, 20] have extended this approach by replacing pose signals with audio signals for talking head generation.
 These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
-Despite these advancements, several critical challenges remain unaddressed.
-First, current approaches suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific requirements; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity.
-Second, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
-Lastly, while existing methods attempt to achieve long-duration portrait animation through sequential frame synthesis, they fail to address the progressive accumulation of residual noise in generated frames.
-This phenomenon inevitably introduces cumulative artifacts (e.g., facial distortions or flickering) that amplify over extended sequences, ultimately degrading output fidelity.
-In this work, we introduce LetsTalk, a linear diffusion transformer for efficient generation of realistic, long-duration talking videos under multimodal conditions.
-Built upon a diffusion transformer as the video generation backbone, LetsTalk replaces conventional text-to-image architectures with an efficient design tailored for audio-driven animation.
-Key to our approach is a deep compression autoencoder that compresses input frames 32×, drastically reducing the number of latent tokens while preserving critical visual details.
-Each transformer block incorporates dedicated temporal and spatial modules, where linear attention (Xie et al., 2024a) ensures efficient modeling of intra-frame and cross-frame coherence without sacrificing computational efficiency.
+However, several critical challenges remain unaddressed, particularly in the context of long-duration video generation.
+As the length of generated videos increases, existing methods struggle to maintain high visual quality, portrait consistency, and temporal coherence.
+Progressive accumulation of residual noise and artifacts (e.g., facial distortions or flickering) becomes increasingly severe over extended sequences, ultimately degrading output fidelity and realism.
+In addition to these long-duration challenges, current approaches also suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific tasks; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity.
+Furthermore, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
+In this work, we introduce LetsTalk, a diffusion transformer framework designed for efficient and high-quality generation of realistic, long-duration talking videos under multimodal conditions.
+A key component for long video generation in LetsTalk is the memory bank mechanism, which enables each generated clip to access information from previous clips, ensuring smooth transitions and contextual continuity.
+To address the discrepancy between training and inference, we propose a noise-regularized training strategy that explicitly simulates inference-time noise accumulation, thereby mitigating error accumulation and sampling artifacts as the video extends.
+[…]
 Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio.
-To better exploit multimodality to guide video generation, we first summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion.
-We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities, and further investigate the underlying factors influencing such adaptability—thereby bridging theoretical gaps in conditional guidance modeling.
-To better guide video generation effectively, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency due to its visual similarity to the video.
-For audio, we employ Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion).
-To enhance long-duration video generation, we incorporate a memory bank mechanism to maintain contextual continuity across sequential video sequences.
-Furthermore, we propose a novel noise-regularized training strategy and a patch drop augmentation to align the memory bank between training and inference.
-By explicitly simulating the residual noise accumulation that occurs during inference (where errors propagate as the video extends), our approach learns to compensate for such artifacts during training for robust long-sequence generation.
-We summarize the contributions of this work as follows:
-  • We present LetsTalk, a linear diffusion transformer for audio-driven portrait animation, which employs efficient linear attention for spatiotemporal modeling and integrates a deep compression autoencoder to achieve 32 $\times$ latent space reduction while preserving visual fidelity;
-  • We systematically analyze three multimodal fusion schemes, characterize their adaptation properties, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality guidance;
-  • We enable long-duration video generation via a memory bank framework with noise-regularized training, aligning memory states between training and inference to mitigate cumulative artifacts in extended sequences;
-  • Our experiments demonstrate our method’s superiority in generating realistic, temporally coherent videos with enhanced diversity, achieving an optimal balance between computational efficiency and output quality.
-2. Related Work
-2.1. Diffusion Models for Video Generation
-Diffusion models have demonstrated exceptional potential across various domains, particularly in image (Ho et al., 2020; Dhariwal and Nichol, 2021) and video generation (Guo et al., 2023; Hu, 2024; Ho et al., 2022b; Zhu et al., 2024).
-A notable example is Stable Diffusion (Rombach et al., 2022), which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets (Schuhmann et al., 2022).
-This approach has been extended to more complex tasks like video generation (Guo et al., 2023; Hu, 2024), where models such as Video Diffusion Models (VDM)(Ho et al., 2022b) apply the diffusion process with space-time factorized training, and ImagenVideo(Ho et al., 2022a) enhances VDM with cascaded diffusion models for high-resolution outputs.
-Further advances, including VideoComposer (Wang et al., 2024b), AnimateDiff (Guo et al., 2023), and VideoCrafter (Chen et al., 2023b), focus on synthesizing videos by conditioning the diffusion process on images and blending image latents.
-A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), which facilitates scaling to larger datasets and model parameters.
-Recent works, such as Latte (Ma et al., 2024) and VDT (Lu et al., 2023), have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch.
+To better exploit multimodality for guiding long-duration video generation, we first summarize three fusion schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion.
+We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities.
+We further investigate the underlying factors influencing such adaptability, thereby bridging theoretical gaps in conditional guidance modeling.
+To effectively guide long-duration video generation, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency throughout extended sequences due to its visual similarity to the video.
+For audio, we adopt Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion) over long video durations.
+We summarize the contributions of this work as follows
+  • We propose LetsTalk, a novel diffusion transformer framework for long-duration talking video generation, which simultaneously achieves high video quality, portrait and temporal consistency, and computational efficiency.
+  • We introduce a memory bank mechanism combined with a noise-regularized training strategy, which explicitly maintains contextual continuity and effectively mitigates error accumulation and sampling artifacts during long video generation.
+  • We design an efficient spatiotemporal modeling architecture, leveraging a deep compression autoencoder and linear attention, which achieves comparable visual fidelity with only 1/8 the parameters of previous methods.
+  • We systematically analyze multimodal fusion schemes and adopt deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio, enabling LetsTalk to achieve state-of-the-art visual quality and temporal coherence in talking video generation.
+II Related Work
+II-A Diffusion Models for Video Generation
+Diffusion models have demonstrated exceptional potential across various domains, particularly in image [11, 12] and video generation [22, 17, 23, 24].
+A notable example is Stable Diffusion [13], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [25].
+This approach has been extended to more complex tasks like video generation [22, 17], where models such as Video Diffusion Models (VDM)[23] apply the diffusion process with space-time factorized training, and ImagenVideo[26] enhances VDM with cascaded diffusion models for high-resolution outputs.
+Further advances, including VideoComposer [27], AnimateDiff [22], and VideoCrafter [28], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents.
+A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [29], which facilitates scaling to larger datasets and model parameters.
+Recent works, such as Latte [15] and VDT [16], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch.
 Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
-2.2. Audio-driven Portrait Image Animation
+II-B Audio-driven Portrait Image Animation
 In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs.
-Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert (Prajwal et al., 2020) achieving high accuracy for specific identities.
-Subsequently, methods such as ATVGnet (Chen et al., 2019) and MakeItTalk (Zhou et al., 2020) introduced facial landmarks and disentangled representations to improve video frame generation.
-The integration of 3D information in models like SadTalker (Zhang et al., 2023a), Genface (Ye et al., 2023), and DiffTalk (Shen et al., 2023) further improved head movements and expressions.
-A significant advancement came with the introduction of diffusion models, including Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), and VividTalk (Sun et al., 2023), which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles.
-Recent innovations like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2024) and AniPortrait (Wei et al., 2024) have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements.
-AniTalker (Liu et al., 2024) introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars.
-LivePortrait (Guo et al., 2024) extends the implicit-keypoint framework, balancing efficiency and controllability.
-Most recently, Hallo (Xu et al., 2024b; Cui et al., 2024), EchoMimic (Chen et al., 2024) and X-Portrait (Xie et al., 2024b) have adopted the latent diffusion framework pioneered in AnimateAnyone (Hu, 2024), adding more control over elements such as lip movements, expressions, and poses.
-2.3. Conditionally Guided Diffusion Model
+Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [7] achieving high accuracy for specific identities.
+Subsequently, methods such as ATVGnet [6] and MakeItTalk [8] introduced facial landmarks and disentangled representations to improve video frame generation.
+The integration of 3D information in models like SadTalker [1], Genface [30], and DiffTalk [31] further improved head movements and expressions.
+A significant advancement came with the introduction of diffusion models, including Diffused Heads [32], DreamTalk [33], and VividTalk [34], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles.
+Recent innovations like VASA-1 [35], EMO [36] and AniPortrait [2] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements.
+AniTalker [37] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars.
+LivePortrait [38] extends the implicit-keypoint framework, balancing efficiency and controllability.
+Most recently, Hallo [18, 20], EchoMimic [19] and X-Portrait [39] have adopted the latent diffusion framework pioneered in AnimateAnyone [17], adding more control over elements such as lip movements, expressions, and poses.
+II-C Conditionally Guided Diffusion Model
 The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research.
-Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms (Ramesh et al., 2022; Avrahami et al., 2022; Brooks et al., 2023; Yang et al., 2024; Khachatryan et al., 2023).
-ControlNet (Zhang et al., 2023b) enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning.
-In video generation, AnimateAnyone (Hu, 2024) introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images.
-Gen-1 (Esser et al., 2023) integrates structural guidance through depth maps and cross-attention, while MCDiff (Chen et al., 2023a), LaMD (Hu et al., 2023), and MotionCtrl (Wang et al., 2024a) focus on motion-guided video generation, enhancing the realism of human actions.
-For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits (Shen et al., 2023; Stypułkowski et al., 2024).
-Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information (Tian et al., 2024; Xu et al., 2024b).
+Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [40, 41, 42, 43, 44].
+ControlNet [14] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning.
+In video generation, AnimateAnyone [17] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images.
+Gen-1 [45] integrates structural guidance through depth maps and cross-attention, while MCDiff [46], LaMD [47], and MotionCtrl [48] focus on motion-guided video generation, enhancing the realism of human actions.
+For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [31, 32].
+Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [36, 18].
 While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio).
 To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
-Figure 2.
-The illustration of our method.
-LetsTalk integrates a deep compression autoencoder that efficiently reduces spatial redundancy while preserving temporal features, and transformer blocks equipped with dual temporal and spatial attention modules, enabling the capture of intra-frame spatial details while establishing temporal correspondences across frames.
-After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding.
-Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding.
-To support long-duration video generation, we introduce a memory bank module to preserve temporal consistency across video sequences.
-Additionally, we devise a novel training strategy to to align memory bank between training and inference.
-By explicitly simulating the residual noise during inference, our approach learns to compensate for such artifacts, ensuring robust long-duration generation.
-3. Method
+III Methodology
+Figure 2: The illustration of our method.
+LetsTalk is designed for robust long-duration talking video generation.
+It integrates a deep compression autoencoder to reduce spatial redundancy while preserving temporal features, and transformer blocks with temporal and spatial attention to capture intra-frame details and long-range temporal correspondences.
+After obtaining portrait and audio embeddings, Symbiotic Fusion is used for the portrait embedding and Direct Fusion for the audio embedding, providing effective multimodal guidance for long video synthesis.
+The portrait embedding is repeated along the frame axis to match the noise embedding for consistent conditioning.
+To support long video generation, we introduce a memory bank module to preserve temporal consistency across extended sequences.
+We also propose a noise-regularized training strategy to align the memory bank between training and inference, enabling stable and high-fidelity long-duration generation.
 We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig.
 2.
-In this section, we first detail our efficient transformer architecture in Sec.
-3.1.
-Next, we outline multimodal fusion schemes in Sec.
-3.2 and describe our strategy for integrating reference portraits with audio in Sec.
-3.3.
-Finally, we introduce our noise-regularized training approach for long-duration video generation Sec.
-3.4.
-3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
-Deep Compression Autoencoder: Building upon latent diffusion models, we employ a deep compression autoencoder (Xie et al., 2024a) that downsamples images by a factor of 32 (vs.
-8 in conventional VAEs), while maintaining reconstruction quality comparable to SDXL’s autoencoder through advanced architectural designs.
-This design is motivated by two key insights: (1) high-resolution video frames contain substantial redundant information that can be safely compressed, and (2) for audio-driven generation, separating compression from denoising enables specialized processing - the autoencoder preserves crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
-The compressed latent feature $\mathcal{F}\in\mathbb{R}^{F\times H\times W\times C}$ is processed through non-overlapping patches, where spatial-temporal positional embeddings (Vaswani, 2017) preserve structural relationships.
-This compact representation enables efficient transformer processing while reducing memory requirements by 4$\times$ compared to conventional approaches.
+This section first introduces the diffusion formulation and transformer backbone in Section III-A.
+We then detail our long-range consistency mechanism, including the memory bank and noise-regularized training, in Section III-C.
+Next, we describe the efficient spatiotemporal transformer backbone in Section III-B.
+Finally, we analyze multimodal fusion schemes and our guidance strategy in Section III-D and Section III-E.
+III-A Preliminaries
+Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters.
+By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
+Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$.
+This is optimized via the variational lower bound of $\log p(x_{0})$:
+$\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
+With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances.
+Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
+$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
+To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
+Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework.
+DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches.
+This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently.
+Different from ViT, DiT incorporates a variational autoencoder (VAE) [49] in the framework, thus runs in the latent space, rather than pixel space.
+Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
+III-B Efficient SpatioTemporal Transformer with Linear Attention
+Deep Compression Autoencoder: We build upon latent diffusion models and employ a deep compression autoencoder [21] that downsamples images by a factor of 32, maintaining reconstruction quality comparable to SDXL’s autoencoder.
+This design leverages the redundancy in high-resolution video frames and decouples compression from denoising, allowing the autoencoder to preserve crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
+Compared to conventional autoencoders, this approach significantly reduces memory and computation requirements while maintaining high fidelity.
+[…]
 Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal.
-While the original DiT (Peebles and Xie, 2023) processes individual frames independently, our video generation framework extends this through interleaved spatiotemporal fusion.
-The naive approach of flattening $\bm{Z}$ into $(F\times P,d)$ tokens with standard attention incurs prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution video generation.
-Our architecture employs a dual-block decomposition strategy that alternates between spatial and temporal modeling, significantly reducing computational complexity from $\mathcal{O}(F^{2}P^{2})$ to $\mathcal{O}(F^{2}+P^{2})$.
-To reduce computational complexity, we replace the traditional softmax attention with ReLU linear attention (Katharopoulos et al., 2020).
-For spatial modeling, the input is first reshaped into $(F,P,d)$ dimensions to compute intra-frame attention using:
-(1) $\bm{Z_{s}}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}$
-where $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$, $\phi(x)=\text{ReLU}(x)$.
-The linear attention mechanism further optimizes this by reusing shared terms across all queries: $\sum_{j=1}^{N}\text{ReLU}(\bm{K}_{j})\bm{V}_{j}^{T}\in\mathbb{R}^{d\times d}$.
-For temporal modeling, the tensor to $(P,F,d)$ dimensions.
-Since $F\ll P$, we preserve the expressive power of standard softmax attention for inter-frame interactions: $\bm{Z_{t}}=\text{Attn}(\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t})=\frac{\bm{V}_{t}(\bm{K}_{t}^{\top}\bm{Q}_{t})}{\sqrt{d}}$.
-Through this hybrid decomposition with linear attention, the overall complexity is reduced to $\mathcal{O}(F^{2}+Pd))$ achieving near-linear efficiency while preserving global receptive fields.
+Unlike the original DiT [29], which processes each frame independently, our framework introduces interleaved spatiotemporal fusion to jointly model spatial and temporal dependencies.
+A naive approach would flatten $\bm{Z}$ into $(F\times P,d)$ tokens and apply standard self-attention, resulting in prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution videos.
+To address this, we adopt a dual-block decomposition strategy that alternates between spatial and temporal attention blocks, significantly reducing the computational cost to $\mathcal{O}(F^{2}+P^{2})$.
+For spatial modeling, we reshape the input to $(F,P,d)$ and apply ReLU-based linear attention [51] within each frame:
+$\bm{Z}_{s}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}},$ (3)
+where $\phi(x)=\text{ReLU}(x)$ and $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$.
+For temporal modeling, we transpose the tensor to $(P,F,d)$ and apply standard softmax attention across frames, allowing us to preserve the expressive power of full attention for inter-frame interactions.
+By interleaving these spatial and temporal attention blocks, our architecture achieves near-linear efficiency while maintaining a global receptive field, enabling scalable and effective modeling of high-resolution, long-duration talking videos.
 Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance.
-Following SANA (Xie et al., 2024a), we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU.
+Following SANA [21], we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU.
 The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality.
-Our experiments demonstrate this architecture achieves superior computational efficiency without compromising visual quality or temporal consistency compared to quadratic-complexity baselines in Fig.
-1.
-Figure 3.
-Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention.
+This enhances local feature capture and compensates for the weaker locality of linear attention.
+Adaptive Integration Following VDT [16], we insert an adaptive fully-connected layer between attention layers to fuse temporal dynamics into spatial representations:
+$\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$ (4)
+where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
+Figure 3: The illustration of the long-duration generation.
+III-C Memory Bank and Noise-Regularized Training for Long-Term Consistency
+In diffusion-based video generation, maintaining temporal alignment and consistency across long sequences is essential.
+However, as the video extends, the accumulation of residual noise and sampling artifacts can threaten temporal coherence and visual quality.
+This is further exacerbated by the mismatch between training and inference: during training, the memory bank is constructed from ground-truth frames, while during inference, it is updated with generated frames.
+Memory Bank: To address these challenges, we introduce a memory bank that stores information from previously generated clips.
+Unlike previous frame-by-frame generation approaches [31, 32], our end-to-end diffusion transformer generates an entire video clip in a single sampling process.
+The memory bank is patchified and temporally transposed, resulting in $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ denotes the number of frames selected from prior clips.
+The memory bank provides prior information to the current generation through Symbiotic Fusion with $\bm{Z_{t}}$, leveraging the scalability of the Transformer architecture.
+This design ensures smooth transitions and temporal consistency across video segments, and is especially advantageous for long-duration video generation.
+Theoretical Motivation: We observed that DDPM-based sampling cannot fully remove noise due to two main factors:
+  • Model prediction error: The predicted noise $\epsilon_{\theta}(x_{t},t)$ only approximates the true noise $\epsilon$, leaving some residual noise after denoising.
+  • Sampling stochasticity: At each reverse step, new Gaussian noise is added:
+    $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)+\sigma_{t}z,\quad z\sim\mathcal{N}(0,I)$ (5)
+    The term $\sigma_{t}z$ introduces inherent randomness.
+Consequently, even after finite-step denoising, diffusion models retain residual noise, which accumulates as the video sequence extends.
+This motivates our noise-regularized memory bank, which differs from previous long video generation methods.
+Noise-Regularized Training Strategy.
+To bridge the gap between training and inference, we propose a noise-regularized training strategy.
+We perturb the memory bank with controlled Gaussian noise, $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$, simulating inference-time noise accumulation.
+Specifically, we randomly sample a diffusion timestep $t$ from the first 50 steps and add the corresponding noise to the memory bank during training.
+Additionally, inspired by Hallo2 [20], we apply patch-drop augmentation by randomly masking 25% of spatial patches in the memory bank, further enhancing model robustness.
+This deliberate corruption enables the model to learn to correct for artifacts and maintain temporal dependencies.
+Our framework thus achieves robust and stable long-duration video synthesis, as demonstrated in Fig.
+3.
+Figure 4: Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention.
 The backbone architecture (left-side blocks) remains consistent across all approaches.
@@
 Symbiotic Self-attention Dynamic Shared Strong
-Table 1. Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
-3.2. Analysis of Multimodal Fusion Schemes
+TABLE I: Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
+III-D Analysis of Multimodal Fusion Schemes
 After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task.
@@
 Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization.
-Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models (Rombach et al., 2022; Nichol et al., 2021).
-In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data (Wang et al., 2024c).
+Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models [13, 52].
+In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data [53].
 Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
 Condition States: The static versus dynamic nature of conditions reflects their constraint strength.
-Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet (Hu, 2024) and ControlNet (Zhang et al., 2023b) that multi-level conditioning enhances generation controllability.
+Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet [17] and ControlNet [14] that multi-level conditioning enhances generation controllability.
 This flexibility improves feature alignment and model performance.
-Parameters Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters.
-Direct Attention, relying solely on cross-attention, also does not share parameters.
-Modality Adaptation: The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks.
+Parameter Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters.
+Direct Fusion, relying solely on cross-attention, also does not share parameters.
+Parameter sharing in Symbiotic Fusion is particularly advantageous for large-scale training, improving efficiency and scalability.
+Modality Adaptation.
+The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks.
 Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios.
 Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment.
-This is exemplified in models such as ControlNet (Zhang et al., 2023b), which employs visual cue-level alignment, and AnimateAnyone (Hu, 2024), which utilizes pixel-level alignment.
+This is exemplified in models such as ControlNet [14], which employs visual cue-level alignment, and AnimateAnyone [17], which utilizes pixel-level alignment.
 The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand.
 Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters.
-This makes it particularly effective for large-scale applications, such as AnimateAnyone (Hu, 2024), where shared parameters ensure robust and efficient multimodal integration.
-The comparative analysis of the three fusion approaches are summarized in Tab.
-1
-3.3. Guidance on Reference Portrait and Audio
-Given the reference image, we require that the character of the generated video is consistent with the reference.
-Since the image modality is close to the video, Symbiotic Fusion proves appropriate.
-Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space.
-Then we temporally repeat the portrait feature $\bm{P}$ with $F$ frames and concatenate it with the latent feature $\bm{Z_{s}}\in\mathbb{R}^{F\times P\times d}$ along the spatial dimension, which can be formulated as: $\bm{Z}=\text{concat}\left(\,\bm{Z_{s}},\,\tilde{\bm{P}}\,\right)$ denotes the temporally extended portrait feature.
-This operation yields a concatenated latent feature $\bm{Z}\in\mathbb{R}^{F\times 2P\times d}$.
-By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames.
-In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers.
-While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
-As for the audio, we require it to reasonably drive the animation, especially the lip motion (Zhang et al., 2023a).
+This makes it particularly effective for large-scale applications, such as AnimateAnyone [17], where shared parameters ensure robust and efficient multimodal integration.
+The comparative analysis of the three fusion approaches is summarized in Table I.
+In summary, we adopt Symbiotic Fusion for portrait features to ensure strong visual consistency, and Direct Fusion for audio features to maintain motion diversity and effective audio-visual alignment.
+III-E Guidance on Reference Portrait and Audio
+[…]
+As for the audio, we require it to reasonably drive the animation, especially the lip motion [1].
 However, certain unconscious actions (e.g.
-blinking and head movements) are almost independent of the audio yet serve to diversity.
-Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity.
+blinking and head movements) are almost independent of the audio yet serve to increase diversity.
+Considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity.
 In a nutshell, Direct Fusion is concise yet effective for fusing the input audio.
-In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow (Xu et al., 2024b) and extract richer semantic information from different layers of the wav2vec encoder (Schneider et al., 2019).
+For the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [18] and extract richer semantic information from different layers of the wav2vec encoder [54].
 Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space.
 Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation.
-The experiments in Sec.
-4.4 verify the efficacy of our mechanism in feature fusion.
-Figure 4. The illustration of the long-duration generation.
-3.4. Long-Duration Generation
-Unlike the previous frame-by-frame generation approach (Shen et al., 2023; Stypułkowski et al., 2024), our end-to-end diffusion transformer generates a video clip in a single sampling process.
-In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences.
-Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible.
-To address this, we introduce a memory bank that stores information from previously generated clips.
-This data undergoes patchification and temporal transpose operations, transforming it into $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ represents the number of frames selected from the prior clips.
-Leveraging the scalability of the Transformer architecture, the memory bank directly provides prior information through Symbiotic Fusion with $\bm{Z_{t}}$.
-However, the inherent distribution discrepancy between training (real data) and inference (synthetic data) poses a fundamental challenge for diffusion models, with error accumulation becoming increasingly severe as video length grows.
-To mitigate this latent risk, we develop a noise-regularized training strategy.
-Specifically, we first establish that diffusion models inherently retain residual noise even after finite-step denoising due to the discrete approximation of continuous noise schedules, which progressively accumulates as the video sequence extends.
-Building on the insight, we introduce a bounded stochastic noiser that perturbs the memory bank contents with controlled Gaussian noise $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$ (see Fig.
-4) into the memory bank during training.
-Inspired by Hallo2 (Cui et al., 2024), we further apply a patch-drop data augmentation strategy to the memory bank.
-Unlike Hallo2 which protects face patches from dropping, we employ completely random dropping to enhance the model’s robustness to the memory bank.
-This deliberate corruption serves two purposes: (1) it faithfully simulates inference-time distribution shifts, and (2) it preserves crucial temporal dependencies.
-The resulting framework demonstrates significantly improved robustness against error propagation in long-duration generation tasks.
-Figure 5. The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
-4. Experiment
-4.1. Experimental Setups
-Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022).
+The experiments in Section IV-F verify the efficacy of our mechanism in feature fusion.
+IV Experiment
+Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
+Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
+IV-A Experimental Setups
+Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [55] and CelebV-HQ [56].
 These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds.
-To ensure high-quality training data, we adopt the preprocess scheme used in (Xu et al., 2024b).
+To ensure high-quality training data, we adopt the preprocess scheme used in [18].
 We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots.
@@
 After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset.
-During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
+During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $512\times 512$ resolution.
 Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID).
-Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs (Karras et al., 2019).
+Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [57].
 SSIM indices measure the structural similarity between ground truth videos and generated videos.
 E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure.
-It extracts facial expression parameters via a reconstruction method (Deng et al., 2019) and computes FID to quantify the gap between generated and ground truth expressions.
-Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) (Ma et al., 2023), SadTalker (138M) (Zhang et al., 2023a), AniPortrait (2.8B) (Wei et al., 2024), Hallo (2.4B) (Xu et al., 2024b), Hallo (2.4B) (Cui et al., 2024) and EchoMimic (2.6B) (Chen et al., 2024).
+It extracts facial expression parameters via a reconstruction method [58] and computes FID to quantify the gap between generated and ground truth expressions.
+Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [33], SadTalker (138M) [1], AniPortrait (2.8B) [2], Hallo (2.4B) [18], Hallo 2 (2.4B) [20] and EchoMimic (2.6B) [19].
 The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
-Figure 6. The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
+Figure 7: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
 Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-DreamTalk 105.434 529.510 0.406 3.466
-SadTalker 102.371 362.634 0.414 2.168
-AniPortrait 78.284 331.117 0.623 2.903
-EchoMimic 55.791 299.343 0.639 2.493
-Hallo 45.754 285.988 0.698 2.235
-Hallo2 44.343 241.628 0.657 2.017
+DreamTalk [33] 105.434 529.510 0.406 3.466
+SadTalker [1] 102.371 362.634 0.414 2.168
+AniPortrait [2] 78.284 331.117 0.623 2.903
+EchoMimic [19] 55.791 299.343 0.639 2.493
+Hallo [18] 45.754 285.988 0.698 2.235
+Hallo2 [20] 44.343 241.628 0.657 2.017
 Ours (Base) 46.240 289.385 0.666 2.587
 Ours (Large) 31.761 219.119 0.632 1.962
-Table 2. The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
-Implementation details Our approach, in contrast to previous methods (Tian et al., 2024; Xu et al., 2024b; Wei et al., 2024) that rely on a two-stage training process, adopts an end-to-end training strategy.
-We set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
-4.2. Quantitative Results
-The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab.
-2 and Tab.
-3, respectively.
-The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment.
-Notably, our approach achieves the lowest FID (31.761) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset.
-These results show that our method generates both realistic and temporally consistent talking-head videos.
-For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
-4.3. Qualitative Results
+TABLE II: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
+Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
+DreamTalk [33] 64.297 293.905 0.448 5.915
+SadTalker [1] 42.017 77.802 0.469 3.372
+AniPortrait [2] 33.408 107.236 0.578 3.722
+EchoMimic [19] 12.932 59.944 0.567 3.671
+Hallo [18] 8.504 46.431 0.580 3.280
+Hallo2 [20] 8.734 41.213 0.543 3.128
+Ours (Base) 9.585 36.159 0.598 3.349
+Ours (Large) 7.610 33.175 0.584 3.038
+TABLE III: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
+IV-B Additional Implementation Details
+Unlike previous methods [36, 18, 2] that use a two-stage training process, our approach adopts an end-to-end training strategy.
+We provide two model sizes: a Base model for simple fine-tuning, and a Large model with better generalization.
+Diffusion We follow the hyperparameters from ADM [12], using a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
+We also use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their method for timestep embedding.
+Training LetsTalk is initialized from a pre-trained Latte [15] model.
+We use gradient clipping for stability and the AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$ and no weight decay, following Latte’s training setup.
+As in previous generative modeling work [29, 59], we maintain an exponential moving average (EMA) of model weights with a decay rate of $0.9999$, and report all results using the EMA model.
+Inference At inference, the model takes a single reference portrait and driving audio as input, and generates a video sequence that animates the reference image according to the audio.
+The generated video has the same length and resolution as during training: $16$ frames at $512\times 512$ resolution.
+IV-C Quantitative Results
+We report comprehensive quantitative comparisons on the HDTF and CelebV-HQ datasets in Table II and Table III, respectively.
+As shown in the tables, our method achieves the best performance on most metrics.
+Specifically, on the HDTF dataset, LetsTalk-L achieves the lowest FID (31.761) and FVD (219.119), outperforming the previous best method Hallo2 by 28.3% and 9.3%, respectively.
+On CelebV-HQ, our method also achieves the lowest FID (7.610) and FVD (33.175).
+These results demonstrate that our approach generates more realistic and temporally consistent talking-head videos.
+All methods generate the first video clip per validation sample and are evaluated at the same resolution.
+IV-D Qualitative Results
 We have visualized the results of different methods on the HDTF dataset (Fig.
-8) and the CelebV-HQ dataset (Fig.
-9).
+5) and the CelebV-HQ dataset (Fig.
+6).
 From the visualization results, it is evident that our method performs exceptionally well on both datasets.
@@
 Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
-Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-DreamTalk 64.297 293.905 0.448 5.915
-SadTalker 42.017 77.802 0.469 3.372
-AniPortrait 33.408 107.236 0.578 3.722
-EchoMimic 12.932 59.944 0.567 3.671
-Hallo 8.504 46.431 0.580 3.280
-Hallo2 8.734 41.213 0.543 3.128
-Ours (Base) 9.585 36.159 0.598 3.349
-Ours (Large) 7.610 33.175 0.584 3.038
-Table 3. The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
-Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-Direct Fusion 52.999 389.01 0.497 3.149
-Siamese Fusion 47.779 286.692 0.651 2.597
-Symbiotic Fusion 46.240 289.385 0.666 2.587
-Table 4. Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
-Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
-Symbiotic Fusion 51.880 403.628 0.471 3.398
-Siamese Fusion 48.041 318.637 0.585 2.706
-Direct Fusion 46.240 289.385 0.666 2.587
-Table 5. Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
-4.4. Ablation Study
+Figure 8: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
+Method #Params Inference Time (ms/frame) FID $\downarrow$
+SadTalker [1] 138M 23 93
+EchoMimic [19] 2.2B 54 55
+AniPortrait [2] 2.6B 74 82
+Hallo [18] 2.4B 115 44
+LetsTalk-B (Ours) 0.3B 27 46
+LetsTalk-L (Ours) 0.7B 40 37
+TABLE IV: Comparison of model size, inference time (ms per frame), and FID on the HDTF dataset.
+IV-E Efficiency Comparison
+To further demonstrate the practical value of our method, we compare model size, inference speed, and FID score with previous state-of-the-art methods.
+As shown in Table IV, LetsTalk-B achieves comparable or better visual quality with significantly fewer parameters and faster inference than recent diffusion-based approaches.
+The larger LetsTalk-L further improves FID while maintaining reasonable efficiency.
+These results highlight the strong trade-off between quality and efficiency achieved by LetsTalk, making it well-suited for real-world applications.
+IV-F Ablation Study
 In this section, we analyze the impact of different fusion schemes as illustrated in Fig.
@@
 The parameters of these transformers are simultaneously updated during training.
+Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
+Direct Fusion 52.999 389.01 0.497 3.149
+Siamese Fusion 47.779 286.692 0.651 2.597
+Symbiotic Fusion 46.240 289.385 0.666 2.587
+TABLE V: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
+Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
+Symbiotic Fusion 51.880 403.628 0.471 3.398
+Siamese Fusion 48.041 318.637 0.585 2.706
+Direct Fusion 46.240 289.385 0.666 2.587
+TABLE VI: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
 Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used.
-We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab.
-4.
+We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Table V.
 Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality.
@@
 Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity.
-Audio-animation alignment refers to specific visual cues (Zhang et al., 2023a) (e.g.
+Audio-animation alignment refers to specific visual cues [1] (e.g.
 facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g.
@@
 Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity.
-Quantitative results in Tab.
-5 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
-Method FID $\downarrow$ FVD $\downarrow$
-Baseline model with Memory Bank 52.243 276.16
-+ Regularized Training (w/o patch drop) 32.746 264.709
-+ Regularized Training (w patch drop) 34.821 251.684
-Table 6.
-Experimental results of long video generation.
-Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
-4.5. Long-Duration Generation
-To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation.
-Our results in Tab.
-6 demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality.
-The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709).
-On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
-5. Conclusion
-We introduced LetsTalk, an efficient linear diffusion transformer for high-fidelity talking head synthesis, tackling core challenges in multimodal fusion and long-term temporal coherence.
-Our approach integrates a deep compression autoencoder for compact latent representation, alongside a systematic exploration of fusion strategies—yielding an optimal framework with Symbiotic Fusion for portrait consistency and Direct Fusion for audio-video synchronization.
-The memory bank mechanism and noise-regularized training further mitigate error accumulation in extended sequences.
-Extensive experiments demonstrate LetsTalk’s state-of-the-art performance, generating lifelike, temporally coherent animations with enhanced diversity, all while maintaining superior efficiency through its optimized architecture.
-Appendix A Preliminaries
-Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters.
-By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
-Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$.
-This is optimized via the variational lower bound of $\log p(x_{0})$:
-(2) $\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$
-With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances.
-Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
-(3) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$
-To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
-Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework.
-DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches.
-This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently.
-Different from ViT, DiT incorporates a variational autoencoder (VAE) (Kingma, 2013) in the framework, thus runs in the latent space, rather than pixel space.
-Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
-Model Patch size Layer num N Hidden size D Paramaters
-LetsTalk-B 1 12 768  [^2] M
-LetsTalk-L 1 24 1024 773
-Table 7. Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
-Appendix B Network Architecture Details
-B.1. Backbone
-In LetsTalk, we borrow the Variational AutoEncoder (VAE) (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022).
-Given the reference portrait image and noisy video, we first encode them using the encoder in VAE.
-Then we patchify and flatten them along the spatial dimension.
-Finally, we use an MLP to project them into the latent space.
-The embeddings in the latent space will be fed into our transformer backbone.
-Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads.
-Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab.
-7.
-Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$.
-Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$.
-Take our base model (i.e.
-LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
-B.2. Adaptive Integration
-Following VDT(Lu et al., 2023), we insert an adaptive full-connected layer between attention layers:
-(4) $\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$
-where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
-B.3. Audio Encoder
-We resample the audio signal to a 16 KHz sample rate.
-For audio feature extraction, we use wav2vec (Schneider et al., 2019), employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation.
-To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips.
-Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information.
-To extract audio representations suitable for transformer tokenization, we designed an audio projection module.
-This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments).
-Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
-B.4. Siamese Transformer
-The Siamese Fusion scheme has been adopted in AnimateAnyone (Hu, 2024), EMO (Tian et al., 2024), Hallo (Xu et al., 2024b), and so on.
-The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise.
-Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network.
-Considering LetsTalk as a Diffusion Transformer, we introduce a ViT (Dosovitskiy, 2020) pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image.
-For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
-Figure 7.
-Portrait animation results without audio guidance for three fusion schemes.
+Quantitative results in Table VI indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
+Figure 9: Portrait animation results without audio guidance for three fusion schemes.
 Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
-B.5. Impact of Audio-guidance on Diversity
-To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 00 for fusion.
+Impact of Audio-guidance on Diversity To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 0 for fusion.
 and compare the diversity of character actions in the generated videos.
@@
 The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
-Appendix C Additional Implementation Details
-C.1. Diffusion
-We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE.
-After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder.
-We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
-Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
-C.2. Training
-We initialize LetsTalk from a pre-trained Latte (Ma et al., 2024) model on the FaceForensics dataset (Rössler et al., 2018).
-Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion.
-For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay.
-The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips.
-Each model is trained for 300k to 500k iterations.
-Following common practices in generative modeling works (Peebles and Xie, 2023; Bao et al., 2023), an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model.
-Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training.
-Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
-C.3. Inference
-During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio.
-The length and resolution of the generated video clip during inference need to remain consistent with those used during training, which in our experiments are $16$ frames and a resolution of $256\times 256$.
-We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
-C.4. Diffusion
-We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE.
-After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder.
-We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$.
-Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
-Appendix D Qualitative Comparisons
-In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk.
-We show more results on the HDTF (Zhang et al., 2021) and Celeb-HQ (Zhu et al., 2022) benchmarks ( Fig.
-8 and Fig.
-9).
-We show the portrait animation results of LetsTalk driven by English (Fig.
-10), Chinese (Fig.
-11), and singing audios (Fig.
-12).
-We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig.
-13).
-Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
-Figure 8. The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
-Figure 9. The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
-Appendix E Visualization
-We provide video visualizations in our attachment.
-Notably, the model used to generate the video is only trained on HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022), which only include English conversation scenarios of real portraits.
-In addition, we use Chinese audio and singing audio to drive real portraits.
-Furthermore, LetsTalk can drive AI-generated portraits [^2] We use Doubao to generate AI-generated portraits in the attached video..
-The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
-Figure 10. English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 11. Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 12. Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
-Figure 13. Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.
+IV-G User Study
+To comprehensively assess the perceptual quality of our approach, we conducted a user study with 12 participants.
+Each participant was shown videos generated by three different methods: Hallo, EchoMimic, and our proposed LetsTalk-L.
+For the evaluation, 20 audio samples were randomly selected, and for each audio, participants were asked to rank the three corresponding videos according to two criteria: realism, which measures how natural and lifelike the video appears, and synchronization, which evaluates how well the lip movements match the audio.
+The results of the user study are summarized in Fig 10.
+The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization.
+As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories.
+These findings indicate that users consistently preferred the videos generated by our method, both in terms of visual realism and the accuracy of lip synchronization with the audio.
+Overall, the user study demonstrates the effectiveness of our approach in producing high-quality, realistic, and well-synchronized talking head videos.
+Figure 10: User study results (%) on (a) realism (left) and (b) synchronization (right).
+Figure 11: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
+IV-H Long-Duration Generation
+To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation.
+Our results in Table VII demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality.
+The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709).
+On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
+[…]
+52.243 276.16
+43.728 263.173
+32.746 264.709
+34.821 251.684
+TABLE VII: Experimental results of long video generation.
+Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
+✓ indicates the component is enabled.
+IV-I Visualization
+We provide video visualizations showcasing long-duration results of up to 2 minutes.
+In these videos, each row corresponds to a different portrait driven by the same audio segment, demonstrating our model’s ability to generate consistent and diverse talking head videos across various identities.
+Notably, our model is trained only on HDTF [55] and CelebV-HQ [56], which feature English conversational scenarios with real portraits.
+To further demonstrate the generalization ability of our approach, we also present results where real portraits are driven by Chinese speech and singing audio, even though such scenarios are not included in the training data.
+V Conclusion
+In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance.
+LetsTalk introduces a memory bank mechanism and noise-regularized training to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation.
+Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference.
+Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion.
+Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios.
+Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.